### v1

# Comprehensive Review and Empirical Evaluation of Causal Discovery Algorithms for Numerical Data

Abstract
Causal analysis has become an essential component in understanding the underlying causes of phenomena across various fields. Despite its significance, the existing literature on causal discovery algorithms is fragmented, with inconsistent methodologies and a lack of comprehensive evaluations. This study addresses these gaps by conducting an exhaustive review and empirical evaluation of causal discovery methods for numerical data, aiming to provide a clearer and more structured understanding of the field. Our research began with a comprehensive literature review spanning over a decade, revealing that existing surveys fall short in covering the vast array of causal discovery advancements. We meticulously analyzed over 200 scholarly articles to identify 24 distinct algorithms. This extensive analysis led to the development of a novel taxonomy tailored to the complexities of causal discovery, categorizing methods into six main types. Addressing the lack of comprehensive evaluations, our study conducts an extensive empirical assessment of more than 20 causal discovery algorithms on synthetic and real-world datasets. We categorize synthetic datasets based on size, linearity, and noise distribution, employing 5 evaluation metrics, and summarized the top-3 algorithm recommendations for different data scenarios. The recommendations have been validated on 2 real-world datasets. Our results highlight the significant impact of dataset characteristics on algorithm performance. Moreover, a metadata extraction strategy was developed to assist users in algorithm selection on unknown datasets. The accuracy of estimating metadata is higher than 80%. Based on these insights, we offer professional and practical recommendations to help users choose the most suitable causal discovery methods for their specific dataset needs.
Keywords: Causal Discovery, Time Series, Independent and Identically Distributed (I.I.D.) Data, Algorithm Evaluation, Survey
1 Introduction
Causality Research, as a dynamically evolving interdisciplinary field, has garnered increasing scholarly attention. This section elaborates on the fundamental concepts, evolution, and classifications of causal analysis, addresses the limitations of previous studies, and introduces the innovative contributions and specific objectives of the current work.
Causal analysis employs a systematic approach to uncovering the underlying causes of phenomena, primarily addressing the question of “Why.” Since Granger’s seminal work in 1969 (Granger, 1969), which introduced a mathematical concept of causality, the field has expanded from philosophy into economics (Imbens, 2004) and other domains such as medicine (Mani and Cooper, 2000), environmental science (Li et al., 2014), and dynamics (Hu et al., 2015). Recently, the rapid advancement of Artificial Intelligence (AI) has opened new avenues for causal analysis. The integration of machine learning has enhanced the precision and efficiency of data processing for causal inference, while causal learning has established a more reliable and trustworthy framework for machine learning (Scholkopf, 2019; Makhlouf et al., 2020). These two disciplines mutually reinforce each other, driving significant advancements in scientific research.
Given that correlation does not imply causation, causal research necessitates a thorough investigation beyond simple association analysis. Pearl’s work has provided a widely accepted framework (Pearl et al., 2000), known as the “Ladder of Causation,” which delineates three stages: Association, Intervention, and Counterfactual. The initial stage, Association, involves observing relationships between variables, yet it is insufficient for identifying confounders or selection bias that may lead to spurious causation (Cheng et al., 2019). The second stage, Intervention, involves controlled experiments to quantify the causal impact of one variable on another. The final stage, Counterfactual analysis, requires a deep understanding of the causal mechanisms underlying the phenomena. Figure 1 shows the causation ladder and corresponding analysis engine (Bareinboim and Pearl, 2016).
Figure 1: Ladder of Causation
Research on causal analysis is primarily categorized into two areas (Nogueira et al., 2022): Causal Inference and Causal Discovery. Causal Inference typically progresses from cause to effect, focusing on the quantitative problems of “Intervention Stage” within the causation ladder framework (Peters et al., 2017). The central concept of inference involves controlled trials (Yao et al., 2021), where both experimental and control groups are observed to determine the effects of interventions. In contrast, Causal Discovery seeks to identify causal relationships from observed outcomes, emphasizing the qualitative problems to learn causal structures (Gelman, 2011). Upon uncovering causal mechanisms, it becomes possible to infer outcomes based on hypothetical scenarios that have not occurred. Despite their inverse logical relationship, causal discovery and causal inference differ significantly in their research methodologies, algorithms, and applications.
In practical scenarios, a significant challenge in causal analysis is managing diverse and complex data types. Standard data formats include time series (Eichler, 2012), cross-sectional, and panel data. Panel data, which has two dimensions (time and samples), combines elements of cross-sectional data, where all sample observations at a specific time point are included, and time series data, where observations of the same sample are recorded at different time points. Since time series data does not adhere to the assumptions of independent and identically distributed (i.i.d.) random variables, it necessitates specialized research on data processing techniques and causal analysis algorithms. Furthermore, due to the importance and widespread use of time series in real-world applications, many researchers have focused extensively on this data type (Assaad et al., 2021; Biswas and Mukherjee, 2024). If a time series can be viewed as a ”list”, then the i.i.d. variable is a ”set” without self-causes. Therefore, causal discovery for i.i.d. data, which differs from time-series causality, has also attracted a lot of research attention (Xie et al., 2019).
This study concentrates on two pivotal elements of causal analysis: (1) Causal Discovery and (2) Time Series and i.i.d. data analysis. An exhaustive investigation was conducted to delve into the methodologies of causal discovery, encompassing principles, algorithmic strategies, and recent advancements in the field.
Currently, the literature lacks a universally applicable algorithm for causal analysis (Edinburgh et al., 2021). Unlike causal inference, the precision of causal discovery heavily relies on the selection of an appropriate causality model. Hence, it is crucial to systematically organize and consolidate existing algorithms to effectively address various scenarios.
Extensive efforts have been made to review and evaluate causal discovery algorithms. To address the challenge of limited benchmark datasets for causal discovery, many researchers have focused on developing data simulators in fields such as industrial systems (Menegozzo et al., 2022), neurology (Tu et al., 2019), and biology (Ma et al., 2023). Some works evaluated causal discovery methods through experiments, but they are not comprehensive as they only consider one type of methods. For example, Sogawa et al. (2010) evaluated the identification accuracy and robustness of linear non-Gaussian methods and its variants. Raghu et al. (2018) compared the performance of four conditional independence-based algorithms on mixed data with latent variables. Ko et al. (2018) summarized estimation of distribution algorithms (EDAs) and compared their performance on four datasets to infer the best ones.
In addition, there are articles that survey various types of methods. Song et al. (2016) and Käding et al. (2021) compared causal discovery methods for bivariates on real-world bench datasets. However, their study did not include analysis of multivariate algorithms. Ombadi et al. (2020) evaluated four causal discovery algorithms on hydrometeorological data, aiming to guide researchers in determining which causal method is most appropriate based on the characteristics of hydrological system. Assaad et al. (2022) not only systematically organized time-series methodologies but also performed thorough evaluations of representative algorithms based on distinct causal structures.
Despite these contributions, most existing literature primarily focuses on theoretical summaries. Even when experiments were conducted, they mainly assessed the impact of causal structures on model performance. Given the often ambiguous causal structure of observational data, it is vital to provide practical and reliable insights from the user’s perspective.
Unlike previous surveys, this project adopts a data-oriented approach, categorizing causal relationships into four types: i.i.d. causality, time-delay causality, instantaneous causality, and causal pairs. To address the research gap that often overlooks the user’s perspective, this study performs a comprehensive analysis starting from the intrinsic characteristics of the data (data assumptions). Motivated by this approach, comparative experiments were conducted, treating data assumptions as experimental factors and employing various algorithms as experimental subjects. These comparisons aim to establish the relationship between algorithms and specific data features within each causality category. Specifically, our pursuit is dedicated to identifying the optimal algorithm, considering various factors, including data size, linearity, stationarity, and noise attributes. By leveraging the extracted data features, users can choose the most appropriate algorithm for their specific needs.
The specific objectives have been enumerated to provide coherent guidance and tasks for this study:
1. Systematic Investigation and Taxonomy of Causal Discovery Methods: Comprehensively collect and categorize methods and algorithms for causal discovery, summarizing the characteristics and applications of these algorithms.
2. Empirical Assessment of Algorithmic Performance through Comparative Experiments: Conduct horizontal comparisons of algorithms on datasets with different features and select appropriate evaluation metrics to assess their performance from multiple perspectives.
3. Recommendation Mechanism for the Optimal Algorithm for Users: Provide meaningful insights and recommendations on the optimal algorithm for specific datasets, offering decision-making suggestions in various application fields.
The primary aim of this research is to develop recommendation algorithms tailored to specific data features through an exhaustive survey of existing methods and rigorous empirical analysis. Key contributions of this study include the creation of a method library featuring 24 algorithms for future experimental use and the execution of 16 comparative analyses to elucidate the relationship between algorithm performance and specific data features. The findings are extended to unknown datasets, highlighting their broader applicability and implications.
The rest of this paper is organized as follows. Section 2 presents a comprehensive literature review, with Section 2.1 covering the fundamental preliminaries of the domain, Section 2.2 discussing assumptions for causality, Section 2.3 reviewing related works, Section 2.4 examining relevant surveys on causal analysis, and identifying research gaps in the current literature. Section 3 details the survey methodology, describing the method design in Section 3.1, the data collection process in Section 3.2, and the analysis of the survey data in Section 3.3. Section 4 focuses on causal discovery algorithms, beginning with Granger-based methods in Section 4.1, followed by condition independence-based methods in Section 4.2, state space dynamic-based methods in Section 4.3, structural equation modeling-based methods in Section 4.4, deep learning-based methods in Section 4.5, and hybrid methods in Section 4.6. Section 5 outlines the empirical study design, introducing the datasets used in Section 5.1, describing the evaluation metrics in Section 5.2, discussing the algorithms in Section 5.3, covering the environment settings in Section 5.4, and presenting the research questions guiding the study in Section 5.5. Section 6 presents the results analysis, addressing the research questions by comparing algorithm performance in Section 6.1, evaluating real-world applicability in Section 6.2, and exploring the generalization to unknown datasets in Section 6.3. Potential threats to validity are addressed in Section 7. Section 8 concludes this study, summarizing the main findings in Section 8.1 and offering recommendations for further work in Section 8.2.
2 Background
Here we will analyze the preliminaries, related survey literature, and research gaps in causal discovery. The section aims to structure the knowledge body of this academic domain systematically.
2.1 Preliminaries
This section presents the fundamental definitions and corresponding notations associated with causal discovery. One needs to state that matrices are denoted by uppercase bold letters, whereas vectors are indicated by lowercase bold letters. Consider the dataset denoted by $X$, which manifests as a $m\times n$ matrix. Here, $x^{n}$ designates the $nth$ variable, while $x$ embodies $m$ observations. This endeavour categorizes observational data into cross-sectional and time-series data, as defined below.
Definition 1 (Cross-sectional Data): Cross-sectional data is a set of observations collected from subjects at one time point.
Note that we mainly focus on a common type of cross-sectional data, namely independent and identically distributed (i.i.d.) data.
Definition 2 (Time Series): Time series is a sequence of data points arranged in temporal order. Given a time series $x^{n}$, an observation at a specific temporal point $t$ is represented as $x^{n}_{t}$.
The concept of time lag is introduced to discern the demarcation between time-delay and instantaneous causality.
Definition 3 (Time Lag): Time Lag $\tau$ refers to the temporal interval between a cause and its effect.
In cases where $\tau>0$, it signifies the occurrence of the cause $\tau$ units of time prior to its effect. This phenomenon is referred to as “time-delay causality”. Nonetheless, circumstances arising from sampling techniques or other factors might occasion an instance wherein $\tau=0$. In such scenarios, the causal latency is deemed insignificant for observation, and this relationship is classified as “instantaneous causality”.
To further depict the causal interdependencies among variables within the dataset, it becomes imperative to introduce the notion of causal graphs.
Definition 4 (Causal Graph): The causal graph $G$ is composed of two subsets: a set of nodes $v$ and a set of edges $\epsilon$. If variable $x^{i}$ is cause of variable $x^{j}$, denoted as $\bm{x^{i}}\rightarrow\bm{x^{j}}$, this relationship manifests as an edge from node $i$ to node $j$ in Directed Acyclic Graphs (DAGs) (Pearl, 1985).
However, when there are hidden variables in the dataset, Maximal Ancestral Graphs (MAGs) (Richardson and Spirtes, 2002) can represent causal relationships. The types of edges in MAGs are as follows:
$\bm{x^{i}}\rightarrow\bm{x^{j}}$: $\bm{x^{i}}$ causes $\bm{x^{j}}$;
$\bm{x^{i}}\leftrightarrow\bm{x^{j}}$: there is a hidden confounder between $\bm{x^{i}}$ and $\bm{x^{j}}$;
$\bm{x^{i}}-\bm{x^{j}}$: there is a hidden effect variable from both $\bm{x^{i}}$ and $\bm{x^{j}}$.
If we do not consider causal directions, a skeleton graph can signify the causal relations between variables. There are only undirected edges in the skeleton graph that represent causal links. For time series, a window graph is common for causal discovery, referring to the causal graph within the maximum time lag window (Assaad et al., 2023). Figure 2 shows examples of these causal graphs.
Figure 2: Causal Graphs
Building upon the precedent definitions, we can elucidate the tasks of causal discovery. Given a dataset $X$, the objectives of causal discovery are the deduction of the causal graph, the quantification of causal strength, and the determination of causal time lags (when the data is time-series). In essence, the goal is to reconstruct the causal mechanism interconnecting the variables. Our focus is centred upon exploring causal graphs, driven by the objective of identifying causal relations within the variable set $v$. To accomplish this, the outcomes of the causal discovery process are encoded using an adjacency matrix $A$.
Definition 5 (Adjacency Matrix): The Adjacency Matrix of causal discovery constitutes a square matrix of dimensions $n\times n$. Specifically, the row vector is set to signify the cause, while the column vector signifies the effect. If the value located at position $(i,j)$ in the matrix is 1, it denotes the presence of a causal link from variable $x^{i}$ to $x^{j}$.
2.2 Assumptions for Causality
When conducting causal discovery tasks, it is crucial to consider certain foundational assumptions. All the methods analyzed in this study are based on at least one of these assumptions. These assumptions help in relating causality to probability densities (Spirtes and Zhang, 2016).
2.2.1 The Causal Markov Assumption
Spirtes and Zhang (2016) argue that in a dataset, all features are independent of their non-effects (nondescendants in the causal graph) conditional on their direct causes (parents in the causal graph). However, it is important to note that the Causal Markov Assumption can be an oversimplification. It assumes that causality is the sole reason for associations between all features, which is not always true, as other kinds of associations exist as well.
For instance, consider the case of ice cream sales and drowning incidents. According to the Causal Markov Assumption, if these two features were present in a dataset and shared an association, it should imply a causal link. However, we know this isn’t true. During summers, both ice cream sales and swimming activities increase. While there might be some conditional dependencies between the two due to common factors, confusing correlation with causality would lead to spurious results.
2.2.2 The Causal Faithfulness Assumption
The Causal Faithfulness Assumption asserts that all and only the true causal links between features are represented in the observed data. This means there are no latent or unobserved confounding variables impacting the observed features and their underlying causal relationships (Spirtes and Zhang, 2016).
For example, consider a dataset of patients in a lung cancer ward, with smoking habits as one of the features and lung cancer as the target variable. To determine whether smoking causes lung cancer, the causal faithfulness assumption holds if there are no unmeasured factors (like genetic vulnerability) linked to both smoking habits and lung cancer but are absent from the observed dataset. If the assumption does not hold, it indicates the presence of confounding factors. For instance, if the dataset lacks information about the patients’ medical history or genetic predisposition, the observed causal link between smoking and lung cancer may be spurious. In such a scenario, smoking habits may appear to cause lung cancer in the data, but the true causal relationship is distorted by latent confounders.
2.2.3 Markov Equivalence Classes (MEC)
Markov Equivalence Classes (MEC) are another important concept in causal discovery. MECs group statistically indistinguishable causal models that make the same predictions about the probability distributions of observed variables (Spirtes and Zhang, 2016). In a given set of variables V, a Markov Equivalence Class represents an aggregation of DAGs that exhibit the same pattern of conditional independence associations among variables in V.
For example, consider two DAGs, G1 and G2. G1 and G2 lie in the same Markov Equivalence Class if, for each pair of variables v1 and v2 in V, v1 and v2 are conditionally independent given all other variables v3 in V in G1 if and only if they are conditionally independent given v3 in G2. Simply put, for G1 and G2 to lie in the same Markov Equivalence Class, they must imply the same set of conditional independence statements among the variables.
To illustrate this further, consider two plausible causal models, M1 and M2, and three variables P, Q, and R:
$M1:P\rightarrow Q\rightarrow R$
$M2:P\leftarrow Q\leftarrow R$
In M1, P causally influences Q, and Q causally influences R. In M2, R causally influences Q, and Q causally influences P. Despite the different causal directions, M1 and M2 lie in the same Markov Equivalence Class because the observed data generated from these models will exhibit the same trends and conditional independence associations (Spirtes and Zhang, 2016).
2.2.4 The Causal Sufficiency Assumption
The Causal Sufficiency Assumption (Spirtes et al., 2001) states that the common causes between any two variables of the variable set $v$ are entirely contained within $v$ itself, thereby excluding the presence of latent confounders. This condition is considered a prerequisite for the efficacy of most causal discovery algorithms.
For example, consider one causal model M and three variables, P, L, and R:
$M:P\leftarrow L\rightarrow R$
If L is an unobserved variable, indicating that L is a hidden confounder of R and Q, then the model M does not satisfy the causal sufficiency assumption. Under this assumption, a directed edge in DAGs represents a causation from a cause to its effect.
2.3 Related Works
In recent years, the topic of causal discovery has garnered significant interest as researchers endeavor to uncover causal relationships from observational and experimental data. To establish a solid foundation for understanding causal discovery, it is crucial to delve into the field’s core concepts and theories. An exhaustive search was conducted within the Web of Science repository to comprehensively encompass the pivotal papers in this domain, curating 301 relevant articles. In order to facilitate a more in-depth examination and exploration of these papers, a network diagram depicting co-cited literature (Figure 3) was generated.
The term “co-citation” means the phenomenon wherein two papers are simultaneously referenced by one or more subsequent papers, thereby establishing interlinked co-citation relationships between the two papers. In Figure 3, the threshold for the minimum number of co-cited articles was set at 15, affording a lucid means to observe literature of close thematic association, thereby furnishing a framework and cluster of causal discovery. We can see from the figure that foundational works by Judea Pearl, Peter Spirtes, and Clark Glymour have laid the groundwork for causal discovery.
Figure 3: Co-citation network
The most prominent node, depicted in red, represents the seminal work of Peter Spirtes (2001). Pearl’s work on causal Bayesian networks and the introduction of causal graphical models significantly advanced the field’s theoretical foundation. These seminal contributions serve as the basis for further research and methodologies. Peter Spirtes introduced nonparametric structural causal models (SCM) (Pearl, 2009) as a formal and intelligible language for articulating causal knowledge and explaining causal notions used in scientific discourse. These include concepts like randomization, intervention, direct and indirect effects, confounding, counterfactuals, and attribution. The structural language’s algebraic component corresponds to the potential-outcome framework, while its graphical component incorporates Wright’s method of path diagrams. When combined, these components provide a robust approach for causal inference, addressing long-standing issues in empirical sciences, such as confounding control, policy evaluation, mediation analysis, and the algorithmization of counterfactuals.
Since the 1990s, conditional independence connections in data have been used to recover the underlying causal structure. One such constraint-based algorithm, developed by Spirtes and Glymour, is the PC algorithm. This algorithm starts with an undirected graph and recursively deletes edges based on conditional independence judgments. The PC algorithm assumes no confounders (unobserved direct common causes of two measured variables) and that the causal information retrieved is asymptotically true. However, the PC algorithm has limitations, especially with high-dimensional biological datasets, where its runtime grows exponentially with the number of variables, making it inefficient for large datasets. Additionally, the order of factors in the input data influences the output, reducing its reliability. Notably, Runge et al. (2019) proposed PCMCI, represented in a blue node in Figure 3, which extends the PC algorithm to time series data.
Another constraint-based algorithm is the FCI algorithm, also developed by Spirtes et al. The FCI algorithm starts with a fully connected undirected graph and removes edges connecting conditionally independent variables to create a causal graph. In the second phase, it orients edges by recognizing the “V” and “Y” structures and then orients the remaining edges using a set of criteria. FCI can produce asymptotically correct results even in the presence of confounders, making it more robust than PC in dealing with latent confounders. These methods can handle a wide range of data distributions and causal links and they have efficient means of conditional independence verification. It should be noted, however, that they may not always convey entire causal information because they yield equivalence classes that represent multiple causal structures satisfying the same conditional independencies (Spirtes and Zhang, 2016).
Score-based algorithms, which facilitate the identification of causal structures by optimizing a well-defined score function, do not entail any confounders. A well-known score-based method is Greedy Equivalence Search (GES) (Chickering, 2002), which directly searches the space of equivalence classes. GES iteratively adds or removes edges to improve the model’s score until no further modifications are possible. Another recent score-based technique is the Greedy Relaxation of the Sparsest Permutation (GRaSP) (Lam et al., 2022) algorithm developed by Lam et al. This algorithm aims to discover the most sparse permutation of variables that fulfills a specific set of causal constraints by greedily adding variables to the permutation while minimizing the impact on the sparsity measure.
Historically, deliberations on the causal relationship amid two factors have concentrated on the standard situation, where the factors included are Gaussian and have a Linear causal relationship. On the other hand, the discrete case has been considered. The Linear Gaussian circumstance, then again, is an exemption in which the causal asymmetry does not uncover itself in the observed data or their joint dispersion. In the event that the causal asymmetry exists, we should conclude that it must have been concealed somehow, either through latent factors or factual predisposition.
For causal discovery from continuous data, numerous techniques based on Functional Causal Models (FCMs) have recently been presented. The effect $Y$ is represented by an FCM as a function of the direct causes $X$ and certain unmeasurable components or noise.
Several variations of the FCM have been demonstrated to be capable of producing distinct causal directions and have found practical uses. The green node attributed to Shimizu (2006), introduces the linear, non-Gaussian, and acyclic model (LiNGAM). $f$ is linear in the LiNGAM and at most one of the noise terms and cause $X$ is Gaussian. The impact Y is further generated by a post nonlinear transformation on the nonlinear effect of the cause $X+$ noise term in the post nonlinear (PNL) causal model.
RCD (Repetitive Causal Discovery) is another causal functional model-based method recently developed by Maeda et. al. (2020) to uncover variable pairs influenced by latent confounders and determine the causal direction between variables unaffected by confounders. Experiments with simulated and real-world data are used to validate RCD’s effectiveness. RCD consists of three major phases. It begins by extracting a collection of ancestors for each variable, which indicate direct or indirect causes. The back-door criterion (Pearl et al., 2000) is used to infer causal directions between observable variables by taking into account the identified shared causes. The procedure is repeated until no additional changes occur. Second, RCD distinguishes direct causes from indirect causes via conditional independence inference. Finally, RCD finds correlated variable pairs with undetermined causal directions to identify pairs of variables impacted by the same latent confounders. As noted by Maeda et al., the experimental results show that RCD is effective at detecting latent confounders, determining causal directions, and detecting variable pairings that are affected by common confounders.
Spirtes and Zhang’s review (2016) comprehensively covers concepts and methodologies for causal inference, including manipulations, causal models, sample predictive modeling, causal predictive modeling, structural equation models, the causal Markov assumption, and the faithfulness assumption. Their review focuses on constraint-based causal structures, structural equation models, and mechanisms for distinguishing cause and effect.
Benozzo et al. (2016) in their recent studies propose a novel approach for causal discovery in time series. The developed technique is a classification- based causal discovery technique by defining a feature space fundamentally based on Granger Causality and by using the Multivariate Auto Regressive (MAR) model as the data generator. Cross Validation was used to assess the performance of the model and the results were quite promising. This makes it evident that this novel approach is a good substitute to the signal processing method for causality detection between two time series. However, as noted by Benozzo et. al., this approach is only limited to three time series at the moment.
Zheng et al. (2018) introduced DAG-NOTEARS in 2018, a method for causal discovery using continuous optimization. This method offers efficient global updates and avoids assumptions about the local graph structure, although it still faces challenges with nonconvex optimization problems. Another approach by Zheng et al., DAG-GNN, uses variational autoencoders with graph neural network architectures to address NP-hardness (Chickering, 1996; Chickering et al., 2004) in causal discovery.
Traditional approaches face NP-hardness due to the combinatorial nature of the task. NP-hardness is a common issue faced by traditional approaches for causal discovery. Zheng et al. proposes another method called DAG Structure Learning with Graph Neural Networks (DAG-GNN) to cater to this issue with the help of a continuous constraint that allows implementation of renowned continuous optimisation techniques by exploring the effectiveness of neural networks as functional approximators to develop a deep generative model. The method uses variational autoencoders parameterized with novel graph neural network architectures. The performance is tested over synthetic and real world datasets and the results achieved are promising.
Another recent advancement in this field was by Zhu et al. (2020) who introduced a novel approach to use Reinforcement Learning to search for the DAG with optimal score by incorporating a predefined score function and penalties for acyclicity. Zhu et al. use the actor-critic algorithm, implementing encoder-decoder models for building the actor. Post experimenting on synthetic and real datasets, the method gives promising results with 30-node graphs. However, dealing with larger graphs is still a challenge, as noted by Zhu et al. Moreover, there’s still scope for development of a more efficient core function and/or implementing other RL algorithms with appropriate early stoppage.
Recently, Wang et al. (2021) proposed a novel approach based on ordering based causal discovery with reinforcement learning. The ordering-based method for causal discovery used in this study is combined with reinforcement learning (RL), where an ordering is created and pruned using variable selection to engender a DAG. The two mechanisms are created using a Markov Decision Process (MDP) formulation along with an encoder-decoder architecture. The proposed approach has been tested on both artificial and real world datasets. Results show positive improvement as compared to other RL-based causal discovery approaches.
In this analysis, we compare and evaluate these novel approaches with conventional ones, considering recent advancements in the domain to provide a comprehensive evaluation of methods for causal discovery.
2.4 Relevant Survey on Causal Analysis
In addition to the basic works and concepts mentioned in the previous section, we also need to investigate causal discovery from a more holistic perspective. Consequently, an exhaustive analysis must be conducted, combined with authoritative surveys over the past five years, to gain a comprehensive understanding of this field.
It is imperative to differentiate the objectives and conceptual frameworks associated with causal discovery and causal inference to further delve into these two notions, as elucidated by Guo et al. (2020). Guo et al. emphasised that causal inference involves tracing the causal path from cause to effect, aiming to understand the impact of manipulating specific variables on others. Within the realm of causal inference, Yao et al. (2021) have conducted an in-depth investigation into the concepts, methods, and applications of causal inference, contributing significantly to the field. Acknowledging the distinct nature of time series data, which differs from i.i.d. data, is essential. This distinction presents unique challenges and considerations in analysing causality.
Additionally, to supplement the understanding of these concepts, Nogueira et al. (2022) have analysed and compared causal discovery and inference using software tools, providing practical examples for testing. Their work contributed to the existing body of knowledge by exploring the practical application and evaluation of different approaches.
Focusing on time series, Moraffah et al. (2021) provided a comprehensive examination of causality for such data, offering insights into the generation of time series data, methodologies employed in the causal analysis, and evaluation metrics used to assess causal relationships. Notably, the study enumerated these evaluation metrics’ specific attributes and characteristics, providing valuable information for researchers in selecting appropriate metrics for their analyses. Regrettably, their research remained confined to theoretical realms and has yet to be realized through practical trial.
Another article on evaluation metrics was proposed by Cheng et al. (2022). The research conducted thoroughly examines the evaluation methods employed in causal analysis. Their analysis referred to a wide range of considerations, including the availability and suitability of software packages, algorithms’ effectiveness, and datasets’ appropriateness for evaluating causal learning algorithms. By investigating these aspects, Cheng et al. provided researchers with valuable guidance for selecting appropriate evaluation methods in causal analysis studies.
Charles K. Assaad et al. (2022) have made a notable contribution to the field of time series causal discovery, and their work serves as a pivotal reference for the research project at hand. In the theory field, they presented a comprehensive framework comprising seven distinct categories for analyzing causal relationships. In the empirical field, they employed ten algorithms to assess these methods’ performance across different causal structures.
More recently, Runge, J. et al. (2023a) comprehensively summarized methods of causal discovery and proposed a Question-Assumptions-Data (QAD) template, embedding causal discovery into Pearl’s causal ladder. They also designed a method selector to match the optimal algorithm to different graph assumptions. However, they did not further discuss parametric assumptions about datasets through experimentation.
Hasan, U. et al. (2023) summarized causal discovery methods for i.i.d. data and time series, and collected source code of relevant algorithms. They also tested and compared the performance of 9 algorithms of i.i.d. data and 7 algorithms of time series on benchmark datasets. However, they did not further test and analyze the influence of data assumptions on algorithms performance, such as dependency functions and noise distributions.
Figure 4: Question-Data-Method-Evaluation Template
Drawing upon the analysis mentioned above, we have amalgamated and organized the principal research directions, which are visually represented in Figure 4. By employing the Question-Data-Method-Evaluation (QDME) template, the review aims to provide a clear and structured overview of the diverse areas and subtopics within the field of causal discovery research, enhancing the organization and coherence of the work.
Based on the aforementioned papers, while the algorithms and evaluations of causal discovery have become relatively comprehensive, several unresolved challenges persist, highlighting gaps in the existing body of knowledge. First, many articles employ outdated taxonomies and lack updates on the latest algorithms. Second, most surveys emphasize theoretical analyses, often neglecting the systematic experiments necessary for quantitative assessment. Third, although some articles have conducted experimental comparisons of algorithms, these studies primarily consider causal structures as their experimental factors, overlooking the characteristics of the data. Consequently, there is currently an absence of evaluation systems that consider the user’s perspective.
3 Survey Methodology
This section will introduce how to collect relevant research resources, including literature, codes, metrics, and datasets. Moreover, the last section briefly explains the analytical technologies we used.
3.1 Method Design
A quantitative research approach is employed to gather and analyze papers from databases systematically. This approach facilitates measuring and exploring trends, methods, datasets, and evaluation in the research domain. Google Scholar was chosen as the literature database due to its extensive coverage of scholarly articles. Information acquisition was initiated by employing a keyword search approach. Primarily, Figure 5 demonstrates the research trend of causal discovery, underpinned by the number of articles published during the preceding two decades.
Figure 5: Published articles in “Causal Discovery” over the past two decades (accumulative)
The graph illustrates an escalating annual trend in research endeavours about causal discovery. This discernible growth can be attributed partly to the robust advancement of AI technologies in contemporary times, which has laid the groundwork for algorithmic developments facilitated by enhanced data processing capabilities.
Furthermore, for preparing the programming underpinnings, our endeavour extends more than literature collection to encompass acquiring requisite algorithmic source code. To this end, we searched for the source code associated with each algorithm on the GitHub platform, with the integration and revision of original algorithm codes as deemed necessary.
To facilitate a comprehensive and impartial investigative analysis, this study employs both Comparative Analysis and Case Study methodologies. Specifically, a series of comparative experiments involving diverse algorithms on distinct artificial datasets are conducted. By analyzing the magnitudes of evaluation metrics, a quantitative assessment of algorithmic performance is executed, yielding overarching insights. Ultimately, to affirm the robustness and effectiveness of our findings, the conclusions are corroborated through case studies such as real datasets.
3.2 Data Collection
Here we executed a keyword-based search focusing on aspects within the tree graph above. The subsequent table illustrates the collection result derived from the published paper from 2004 to 2023, as sourced from Google Scholar.
To prevent repetitive retrieval and collection of papers, “Causal Discovery” is used as the basic keyword, with previously retrieved terms excluded when searching for new supplementary terms. It should be noted that we also searched for keywords related to “causal inference” to ensure comprehensive collection of papers within cross fields. We have identified 220 articles with the highest correlation from over 3000 related articles and have summarized the datasets, algorithms, and evaluation metrics for causal discovery. The terms “Hits,” “Title,” and “Body” in Table 1 refer to the number of papers returned by the search, the relevance of the titles to the desired content, and the number of papers that remain after title and body framing, respectively.
Table 1: Paper Collection Result
Keywords Hits Title Body
‘Causal Discovery” + [Survey |||| Overview |||| Review] 980 15 14
“Causal Discovery” + [Bench |||| Benchmark] 347 7 7
“Causal Discovery” + Dataset 619 16 16
“Causal Discovery” + [Evaluation |||| Comparison] 240 18 16
“Causal Discovery” + [Method |||| Algorithm |||| Approach] 410 138 137
“Causal Inference” + [Time Series |||| Cross-sectional |||| I.I.D.] 995 89 30
Overall - 283 220
A selection of widely employed datasets for causal discovery has been identified by synthesising diverse review articles. These datasets encompass both real-world instances and artificial constructs. To streamline ensuing experimental processes, a classification framework has been devised to categorise these datasets into four types of causality relationships. Additionally, the sources of these datasets are also categorized into real-world and synthetic datasets. Specifically, CausalWorld (Ahmed et al., 2020), SynTReN (Van den Bulcke et al., 2006), LUCAS (Guyon et al., 2011), ALARM (Lauritzen and Spiegelhalter, 1988) and ASIA (Lauritzen and Spiegelhalter, 1988) are synthetic datasets, while Tubingen (Mooij et al., 2016), ADNI (Petersen et al., 2010), AntiCD3/CD28 (Sachs et al., 2005), Abalone (Asuncion and Newman, 2007), fMRI (Smith et al., 2011), Causality 4 Climate (Runge et al., 2020), Traffic Prediction (Pan et al., 2018), OHDNOAA (Jangyodsuk et al., 2014), Temperature Ozone (Gong et al., 2017), Sachs (Sachs et al., 2005) and CHILD (Spiegelhalter et al., 1993) are real-world datasets. This categorisation is illustrated in Table 2.
Table 2: Datasets Collection Result
Causality Type Dataset Year Area Source
Pairwise Cause Effect Pairs Challenge (Statnikov et al., 2013) 2013 Various Real-world
Tubingen (Mooij et al., 2016) 2016 Various Real-world
CE-Gauss (Mooij et al., 2016) 2016 - Synthetic
CE-Multi||||Net (Goudet et al., 2018) 2018 - Synthetic
Instantaneous SynTReN (Van den Bulcke et al., 2006) 2006 Biology Synthetic
FLUXNET (Pastorello et al., 2020) 2020 Biogeoscience Synthetic
causaLens (Lawrence et al., 2020) 2021 Various Synthetic
Time-delay fMRI (Smith et al., 2011) 2011 Neuroscience Real-world
FinanceCPT (Kleinberg, 2013) 2012 Economics Synthetic
OHDNOAA (Jangyodsuk et al., 2014) 2014 Hydrologic Real-world
Traffic Prediction (Pan et al., 2018) 2018 Traffic Real-world
Causality 4 Climate (Runge et al., 2020) 2020 Climate Real-world
I.I.D. Data Sachs (Sachs et al., 2005) 2005 Biology Real-world
LUCAS (Guyon et al., 2011) 2011 Medical Synthetic
ALARM (Beinlich et al., 1989) 1989 Belief Networks Semi-synthetic
CHILD (Spiegelhalter et al., 1993) 1993 Medical Real-world
ASIA (Lauritzen and Spiegelhalter, 1988) 1988 Medical Synthetic
Auto-mpg (Driessens and Džeroski, 2005) 2005 Engineering Real-world
Likewise, we have identified performance metrics for assessing causal discovery techniques. These metrics are divided into two overarching families: graph-based metrics (Peters and Bühlmann, 2015) and classification-based metrics, as expounded in Table 3. It is pertinent to highlight that nearly all metric computations necessitate the availability of both estimated DAGs and ground truth DAGs. Therefore, the prudent selection of datasets with well-established ground truth becomes imperative.
Table 3: Metrics Collection Result
Measure Type Metric Notions
Graph distance-based measure Structural Hamming Distance (SHD) Calculate the difference between two (binary) adjacency matrices: each edge that is missing or not in the target graph is counted as an error.
Frobenius Norm Compare the similarity between real matrices and estimation matrices.
Structural Intervention Distance (SID) Estimate the count of erroneously deduced intervention distributions.
Classification-based measure Precision The quotient of true positives (TP) divided by the sum of TP and false positives (FP).
Recall The proportion of TP in relation to the summation of TP and false negatives (FN).
F1 Score The harmonic mean of precision and recall of the learned structure as compared to true causal structure.
FPR The ratio of the edges that are present in the predicted graph but not present in the ground-truth graph.
TPR The ratio of the common edges between the ground-truth and predicted causal graphs over the number of edges in ground-truth graph.
MSE The sum of square of difference between the predicted and the ground-truth causal graphs divided by the total number of nodes.
Area under ROC Curve (AUROC) Area under ROC curve is the area under the curve of recall versus FPR at different thresholds.
3.3 Survey Analysis
Utilizing a comparative analysis framework, we systematically process the performance metrics. We selected a reference algorithm exhibiting superior performance to enhance result lucidity and then calculated residuals for other algorithms compared with the reference. The overall distribution of these residuals is visually represented through the violin plots.
Nevertheless, two specific algorithms may manifest insignificant disparities between their metric values. Hence, we introduce a significance assessment mechanism to pursue a more methodical treatment of data relationships. Given the constraint that each data size consists of merely five datasets, we opt for the non-parametric Mann-Whitney U test (Mann and Whitney, 1947). This approach circumvents the necessity to assume normality in data distribution. Significance is determined based on a p-value threshold of 0.05; when under this threshold, it signals a notable divergence in the performance of the two algorithms.
In addition, a ranking table is formulated to delineate the evaluation outcomes for enhancing the discernibility of inter-algorithm performance disparities. This table is ordered in a descending manner upon the average values. Meanwhile, we calculate the standard deviation for each algorithm’s metric values; reduced standard deviation signifies less dependence of algorithmic performance on data size, which means enhanced stability of algorithms.
4 Causal Discovery Algorithms
Various research articles present diverse taxonomy for causal discovery, yet a universally accepted classification structure does not currently exist. Specifically, in recent years, the rapid advancements in this domain have resulted in the incompleteness and obsolescence of numerous surveys. It is essential to collect and analyse the taxonomy methodologies proposed in papers critically to establish a systematic categorisation of existing methods. Simultaneously, the devised structure should encompass as many algorithms as possible. Thus, this project diligently compiles and summarises the prevailing methods, effectively partitioning causal discovery into six fundamental categories, as shown in Figure 6: Granger-Based, Conditional Independence-Based, State Space Dynamics-Based, Structural Equation Modelling-Based, Deep Learning-Based, and Hybrid Method.
{forest}
[…]
Causal
Discovery, align=center [Granger-Based Methods, align=center [Multivariate Granger Analysis Method (Arize, 1993)] [Extended Granger Causality (Chen et al., 2004)] [Kernel Granger Method (Liao et al., 2009)] [Copula Granger Method (Hu and Liang, 2014)] ] [Condition
Independence-Based
Methods, align=center [Information Theoretic-Based Approach - oCSE (Sun et al., 2015) ] [Causal Network-Based
Approach, align=center, s sep=2pt, l=0.9cm [Constraint-Based, [Peter-Clark (PC) (Kalisch and Bühlman, 2007)] [CD-NOD (Huang et al., 2020)] [PCMCI (Runge et al., 2019)] [Fast Causal Inference (FCI) (Entner and Hoyer, 2010b, a)] ] [Score-Based, [Greedy Equivalence Search (GES) (Chickering, 2002)] [ES (Yuan and Malone, 2013)] [GRaSP (Lam et al., 2022)] [DYNOTEARS (Pamfil et al., 2020)] ] ] ] [State Space
Dynamic-Based
Methods, align=center, [CCM (Sugihara et al., 2012)] [Cross Map Smoothness (CMS) (Ma et al., 2014)] [Inner Composition Alignment (IOTA) (Hempel et al., 2011)] [Pairwise Asymmetric Inference (PAI) (McCracken and Weigel, 2014)] ] [Structural Equation
Modelling-Based
Methods, align=center [Structural Equation Modeling (SEM) (Shimizu et al., 2006)] [LiNGAM-Based, [ICA-LiNGAM (Zhang et al., 2021a)] [Direct LiNGAM (Shimizu et al., 2011)] [VARLiNGAM (Hyvärinen et al., 2010)] [RCD (Maeda and Shimizu, 2020)] [CAM-UV (Maeda and Shimizu, 2021)] ] [Additive Noise Models (ANM) (Hoyer et al., 2008), [TiMINo (Peters et al., 2013)] ] [PNL (Zhang and Hyvarinen, 2012)] [DAGs with NO TEARS (Zheng et al., 2018)] [GOLEM (Ng et al., 2020)] ] [Deep Learning-Based
Methods, align=center [Causal Generative Neural Networks (CGNN) (Goudet et al., 2018)] [DAG-GNN (Zhang et al., 2021a)] [Temporal Causal Discovery Framework (TCDF) (Nauta et al., 2019a)] [Amortized Causal Discovery (ACD) (Löwe et al., 2022)] [Ordering-Based Causal Discovery with Reinforcement Learning (Wang et al., 2021)] [GraNDAG (Lachapelle et al., 2019)] ] [Hybrid Methods, align=center [ARMA-LiNGAM (Kawahara et al., 2011)] [Scalable Causation Discovery Algorithm (SCDA) (Raghu et al., 2018)] [PSDR-TE (Mao and Shang, 2017)] [Information Geometric Causal Inference (IGCI) (Janzing et al., 2012)] [Neural Granger Causality (NGC) (Tank et al., 2021b)] ] ]
Figure 6: Taxonomy for Causal Discovery (The underlined algorithm is only applicable to Time-Series)
4.1 Granger Based Method
Granger causality (GC) is one of the pioneering measurement methods for analysing time series data. Over several decades, undergoing refinement and evolution, it still maintains an irreplaceable position in the contemporary landscape of causal discovery. The core premise of Granger causality postulates that future events do not affect the present or past, while past events potentially impact both the present and the future. When the historical information of variables $x$ and $y$ are included, leading to better predictions for variable $y$ than predictions based solely on the information of $y$, it signifies that variable $x$ is considered the Granger cause of variable $y$. In mathematical notation, the given statement can be expressed as follows (McCracken, 2016):
$P(y_{n+1}\in\bm{A}\mid\bm{\Omega_{n}})\neq P(y_{n+1}\in\bm{A}\mid\bm{\Omega_{n}}-\bm{x})$ (1)
Equation 1 illustrates $x$ Granger causes $y$, wherein the variable $x$ and $y$ represents two discrete time series, and the subscript $n$ corresponds to the time point $t$. The all-encompassing set of information available at all points $t\leq n$ is symbolically denoted as $\Omega_{n}$. To ascertain the Granger relationship based on the aforementioned formula, the Vector Autoregressive (VAR) model (Lütkepohl, 2005) stands as the prevailing technique, built upon the premise of data stationarity and equipped to forecast variable values.
Our attention is directed towards several primary algorithms based on GC. Arize et al. (1993) put forth the Multivariate Granger (MVGC) analysis method to overcome the limitations of Pairwise Granger (PWGC), which can only deal with bivariate data. However, ensuring linearity in real-world datasets can present a significant challenge. To address this issue, Chen et al. (2004) introduced an approach known as Extended Granger Causality (EGC), specifically designed to handle nonlinear data. An alternative model catering to nonlinear time series is the Kernel Granger (KGC) method (Marinazzo et al., 2008; Liao et al., 2009; Marinazzo et al., 2011), showcasing notable attributes such as high accuracy and flexibility. When confronted with continuous time series data, Hu et al. (2014) proposed the Copula Granger method, which is capable of uncovering nonlinear and higher-order causal relationships (Kim et al., 2020; Jang et al., 2022).
4.2 Condition Independence Based Method
The conditional independence-based method exhibits a close association with probability. By quantifying the mutual information (Runge, 2018) between variables, this approach enables the determination of causal relations and causal strength. A fundamental concept in this context is the transfer entropy (Schreiber, 2000), which is defined as follows:
$T_{\bm{x}\rightarrow\bm{y}}=\sum P(y_{n+1},y_{n}^{(k)},x_{n}^{(l)})\log\frac{P(y_{n+1}\mid y_{n}^{(k)},x_{n}^{(l)})}{P(y_{n+1}\mid y_{n}^{(k)})}$ (2)
In contrast to Shannon entropy, the transfer entropy is computed by the Kullback entropy (Kullback, 1997). In this equation, $x_{n}$ represents the value of variable $x$ at the $n$th time point, and likewise for the variable $y$, with the superscript indicating the time delay length. When $T_{\bm{x}\rightarrow\bm{y}}-T_{\bm{y}\rightarrow\bm{x}}>0$, it can be inferred that variable $x$ is the cause of variable $y$; conversely, variable $y$ is the cause of variable $x$.
Causal discovery algorithms based on conditional independence can be categorized into two distinct groups. The first category is the information-theoretic-based approach, with Optimal Causation Entropy (oCSE) (Sun et al., 2014, 2015) being a representative algorithm. oCSE is a two-step discovery algorithm explicitly designed for short time series data. The second approach is causal network-based, where the optimal causal graph is determined through statistical testing. The Peter-Clark (PC) algorithm (Kalisch and Bühlman, 2007) has gained widespread adoption and has proven effective in analyzing high-dimensional time series using causal graphs. Recognizing the potential interference of latent confounders in causal detection, corresponding approaches have been developed. A notable example is the Fast Causal Inference (FCI) algorithm (Zhang, 2008; Entner and Hoyer, 2010b; Spirtes et al., 2013), a classical method that explicitly accounts for unobserved confounders. To further enhance control over false positive rates, Runge et al. introduced the PCMCI method and its variants (Runge et al., 2019; Runge, 2020; Gerhardus and Runge, 2020) by incorporating the MCI test into the PC algorithm. PCMCI is an improvement of PC in time series, which can detect contemporaneous and time-delay effects.
CDS (Fonollosa, 2019) was proposed to detect the causal direction of bivariates. This method measures the statistical characteristics of the joint distribution of marginal variance data after conditioning the bins. This algorithm has been proven to be robust as it has a high AUC in ChaLearn causal pair challenges.
CD-NOD (Constraint-based causal Discovery from heterogeneous/NOnstationary Data) (Huang et al., 2020) is another framework designed to discover causal relationships in data where generating processes change over time or across domains. It detects changing local mechanisms, recovers causal structures, and estimates the driving force behind nonstationarity. This nonparametric method leverages data heterogeneity and connects nonstationarity with soft interventions, demonstrating efficacy on synthetic and real-world datasets like task-fMRI and stock market data.
Here we introduce classic score-based approaches. Compared to the PC algorithm, the Greedy Equivalence Search (GES) algorithm, proposed by Chickering et al. (2002), shows enhanced robustness when dealing with nonstationary data. GRaSP (Greedy Relaxations of Sparsest Permutation) (Lam et al., 2022; Andrews et al., 2023) is designed to efficiently identify DAGs representing causal structures from observational data. It builds on permutation-based reasoning and introduces a novel operation called “tuck” to relax the assumptions required by previous methods like TSP (Triangle Sparsest Permutation) and ESP (Edge Sparsest Permutation). GRaSP consists of three tiers: GRaSP0, GRaSP1, and GRaSP2, each progressively weakening the assumptions and increasing the ability to recover sparser permutations. GRaSP2, the most relaxed form, outperforms several state-of-the-art algorithms in simulations, demonstrating scalability and accuracy for dense graphs and those with over 100 variables. To address causal discovery problems in temporal series, Pamfil et al. (2020) introduced DYNOTEARS, a novel Bayesian network learning algorithm that utilizes score constraints to ascertain the edges within the causal structure graph. DYNOTEARS is capable of effectively handling both instantaneous and delayed causality, making it a versatile tool for causal inference in time series data.
4.3 State Space Dynamic Based Method
The state space dynamics-based method can be regarded as a complementary approach to address a category of data not encompassed by GC. This method investigates the causality of variables within weakly coupled dynamic systems, significantly enhancing the causal discovery capability in ecological, dynamics, and other relevant domains. This method draws inspiration from the Takens theorem (Takens, 1981) and computes the bidirectional cross-correlation between two variables to establish a cross-mapping. To be specific, variable $x$ causes $y$ when $C_{\bm{xy}}>C_{\bm{yx}}$ is satisfied, wherein $C_{\bm{xy}}$ and $C_{\bm{yx}}$ represent the Convergent Cross-Mapping (CCM) correlations from $x$ to $y$ and from $y$ to $x$, respectively. The calculation formula for CCM correlation is as follows, where $\rho$ means the Pearson correlation coefficient.
$C_{\bm{xy}}=[\rho(\bm{y},\bm{y}|\tilde{\bm{x}})]^{2}$ (3)
We assume a classic example to explain cross-mapping further, using variable $x$ to construct shadow manifold $M_{\bm{x}}$, and $\bm{y}$ to construct $M_{\bm{y}}$. If $x$ leads to $y$, then using the neighbouring points of a certain point in $M_{\bm{y}}$ should be able to identify better the neighbouring points of the corresponding point in $M_{\bm{x}}$. Supposing a delay of 1 and the shadow manifold graphs from two directions are shown in Figure 7, where $x$ causes $y$.
(a) $M_{x}$ to $M_{y}$ cross mapping
(b) $M_{y}$ to $M_{x}$ cross mapping
Figure 7: Cross mapping manifold (lag = 1)
Sugihara et al. (2012) introduced the concept of CCM to infer causality. This method proves advantageous for non-separable and weakly connected dynamic systems. An essential feature of CCM is convergence, which means that the longer the time series used (with a larger sample size), the smaller the estimated error of the obtained cross-mapping. To overcome this limitation that necessitates long time series data, Ma et al. (2014) developed Cross Map Smoothness (CMS), specifically designed to handle varying data sizes, particularly short time series. Additionally, Inner Composition Alignment (IOTA) (Hempel et al., 2011; Wang et al., 2014) offers an alternative technique for short-time series analysis. McCracken et al. (2014) proposed Pairwise Asymmetric Inference (PAI) as an exploratory tool for analysing high-dimensional dynamic systems. PAI aids in uncovering causal relationships within complex systems by evaluating asymmetries in pairwise interactions.
4.4 Structural Equation Modelling Based Method
The three methods above are well-suited for identifying time delay causality but may not necessarily be adept at detecting instantaneous causality. On the other hand, the Structural Equation Modeling (SEM) based method represents a significant advancement and revolution in the realm of instantaneous causal discovery. This method ascertains the edges of DAG by establishing a structural equation to solve the coefficient matrix. The most basic form of the structural equation (Shimizu et al., 2006) is as follows:
$\mathbf{X}=\mathbf{BX}+\mathbf{E}$ (4)
The matrix $\mathbf{B}$ is referred to as the coefficient matrix, with its row and column representing the two dimensions of cause and effect. Upon determining $\mathbf{B}$, the causal relation can be discerned. Additionally, the matrix $\mathbf{E}$ represents the noise matrix in the model, usually non-Gaussian noise.
The Linear Non-Gaussian Acyclic Model (LiNGAM) provides a directed acyclic graph that reveals instantaneous causal relations between variables. LiNGAM assumes linear and non-Gaussian independent noise about the data generation method of the system, solving these using Independent Component Analysis (ICA) (Lee and Lee, 1998; Naik and Kumar, 2011). Since ICA algorithms typically employ FastICA and gradient-based algorithms, they may converge to local rather than global optima. To address these issues, Shimizu et al. (2011) proposed the Direct LiNGAM algorithm. Compared to ICA-LiNGAM, Direct LiNGAM produces more stable and reliable results, though it has some drawbacks. One drawback is its slower computational efficiency compared to ICA-LiNGAM. Additionally, its assumptions are relatively strict; in real-world scenarios, data generation mechanisms are often nonlinear and do not conform to its assumptions.
To address the gap in the LiNGAM algorithm family concerning delayed causality, Hyvärinen et al. (2010) designed the VARLiNGAM algorithm. VARLiNGAM operates in two steps: first, it predicts time lag effects using a Vector Autoregression (VAR) model; second, it estimates instantaneous causality by applying LiNGAM.
Building upon the insights from LiNGAM, Hoyer et al. (2008) introduced Additive Noise Models (ANM) to detect nonlinear time series, emphasizing that nonlinearities offer valuable identification power. One shortcoming of this algorithm is its high computational cost, as it involves determining the direction between pairs of variables one-on-one, making the algorithm pairwise causality.
The Post-Nonlinear (PNL) causal model (Zhang and Hyvarinen, 2012; Zhang et al., 2015; Uemura et al., 2022) addresses the complexities of nonlinear effects, inner noise, and measurement distortions in observed variables for causal discovery. Representing each variable as a function of its direct causes, an independent disturbance, and a post-nonlinear distortion, PNL can distinguish between causes and effects, especially in non-Gaussian scenarios. The model’s identifiability has been extensively studied, revealing that it can generally identify causal directions except in specific conditions. Empirical results demonstrate its efficacy in various real-world data sets, making it a robust tool for causal inference in complex systems.
Peters et al. (2013) proposed Time Series Models with Independent Noise (TiMINo) to capture both lagged and instantaneous effects. This model is based on nonlinear independence tests and can perform well even when the dataset does not satisfy the causal sufficiency assumption.
DAGs with NO TEARS (Zheng et al., 2018) is a causal discovery method that uses continuous optimization schemes to learn the structure of Directed Acyclic Graphs (DAGs) from observational data. The name “NO TEARS” stands for “Nonlinear Optimal Transformations for Efficient and Accurate Recovery of Structure,” emphasizing its focus on tackling issues related to learning nonlinear causal relationships. The algorithm transforms the data to make it linear or Gaussian and learns a Structural Equation Model (SEM) linking features in a causal graph. It optimizes the SEM with a focus on fit and sparsity using methods like gradient descent or ADMM. Soft thresholding promotes sparsity by zeroing out edges. The process iterates until convergence criteria are met, then returns a DAG with NO TEARS, representing causal relationships.
GOLEM, introduced by Ng et al. (2020), is a continuous likelihood-based method for causal discovery. It uses a score-based approach with soft sparsity and DAG constraints to maximize the data probability of a linear Gaussian model. GOLEM employs two objective functions to account for noise variances and uses an l1 penalty for complexity. It formulates an unconstrained optimization problem, ensuring the graph remains a DAG under reasonable assumptions. The algorithm utilizes gradient-based optimization methods, with Adam optimizer and GPU acceleration. A post-processing step removes low-weight edges to enhance performance. GOLEM effectively restores DAG structures while managing soft constraints.
Repetitive Causal Discovery (RCD) (Maeda and Shimizu, 2020; Maeda, 2022) is a method for identifying causal structures in data affected by latent confounders. It repeatedly infers causal directions between small sets of observed variables, determining if relationships are influenced by latent confounders. The resulting causal graph uses bi-directed arrows to indicate variables sharing the same latent confounders and directed arrows for causal directions between variables not affected by the same latent confounder. Experimental validation with simulated and real-world data shows that RCD effectively identifies latent confounders and causal directions.
Causal Additive Models with Unobserved Variables (CAM-UV) (Maeda and Shimizu, 2021) handle causal discovery in the presence of unobserved variables, particularly for nonlinear causal relationships. This model extends causal additive models by accounting for both unobserved common causes and intermediate variables. CAM-UV identifies all theoretically possible causal relationships without bias from unobserved variables, avoiding incorrect inferences. Empirical results from artificial and simulated fMRI data confirm CAM-UV’s effectiveness in inferring causal structures despite the presence of unobserved variables.
4.5 Deep Learning Based Method
Deep learning-based methods have emerged as powerful tools in causal discovery, closely connected with machine learning. These methods offer significant technical advantages, particularly in processing vast amounts of data. Notably, deep learning-based causal algorithms can better infer hidden variables using network information.
For instance, Goudet et al. (2018) designed Causal Generative Neural Networks (CGNN) to address the challenges posed by latent variables in causal analysis. CGNN is an algorithm that infers the optimal causal direction on a causal skeleton diagram, which belongs to pairwise causality. Through testing on both artificial and real-world datasets, CGNN has demonstrated advanced performance in handling potential confounders.
DAG-GNN, developed by Yu et al. (2019), combines Graph Neural Networks (GNNs) with a score-based approach to learn DAGs from data. It uses GNNs for node embeddings to model feature dependencies and detect causal relationships. The method starts by embedding features using GNNs, then defines a score to evaluate the causal structure. It formulates an optimization problem to maximize this score using gradient-based techniques like stochastic gradient descent or Adam. Causal constraints ensure acyclicity. The data is split into training and validation sets to optimize the score and train the model. After achieving optimal performance on the validation set, the algorithm returns a DAG representing causal relationships.
Another noteworthy algorithm is the Temporal Causal Discovery Framework (TCDF) (Nauta et al., 2019a), which effectively handles both latent and instantaneous causal effects. TCDF adopts an attention mechanism in convolutional neural networks. The attention coefficients of different variables, learned by the network, can be interpreted as the degree of correlation between variables. If the attention coefficient is below a certain threshold, it indicates no causal relationship between the two variables.
GraNDAG (Lachapelle et al., 2019) is a novel score-based approach for learning DAGs from observational data. It adapts a recent continuous constrained optimization formulation to accommodate nonlinear relationships between variables using neural networks. This method effectively models complex interactions and avoids the combinatorial nature of the problem. By comparing GraNDAG to existing continuous optimization methods and nonlinear greedy search methods, it has been demonstrated that GraNDAG outperforms current continuous methods on most tasks and remains competitive with existing greedy search methods on important causal inference metrics.
Ordering-Based Causal Discovery with Reinforcement Learning (CORL), developed by Wang et al. (2021), combines ordering-based causal discovery with reinforcement learning techniques to learn causal relationships by generating and refining an ordering of variables. The algorithm treats the problem as a sequential decision-making task, where a reinforcement learning agent arranges variables to approximate true causal relationships. A reward function provides feedback, incentivizing correct orderings and penalizing incorrect ones. The task is formulated as a Markov Decision Process (MDP), with states representing the current ordering and actions selecting the next variable. The agent is trained using reinforcement learning algorithms like Q-learning or Proximal Policy Optimization to optimize long-term rewards. Balancing exploration and exploitation is crucial. A post-processing step, such as local search, refines the ordering to improve accuracy. The algorithm ultimately returns an optimal causal graph representing the relationships between features.
More recently, Löwe et al. (2022) introduced the Amortized Causal Discovery (ACD) algorithm for time series, which is effective with small data sample sizes and demonstrates efficacy in dynamic systems. This model utilizes shared information between dynamic system variables to identify confounders in additive noise datasets.
4.6 Hybrid Method
Hybrid methods combine two or more algorithms to complement and optimize each other, enhancing the ability to discover causality. These methods leverage the strengths of different approaches to address their individual limitations and improve overall performance.
One illustrative hybrid approach is the Autoregressive Moving Average - Linear Non-Gaussian Acyclic Model (ARMA-LiNGAM) (Kawahara et al., 2011), which combines Granger Causality (GC) and Structural Equation Modeling (SEM). This composite method resolves both instantaneous and delayed causality by leveraging the attributes of LiNGAM and ARMA models. ARMA-LiNGAM’s integration allows for a more comprehensive analysis of time series data, accommodating both immediate and lagged effects. Incorporating deep learning techniques with GC models, Neural Granger Causality (NGC) (Tank et al., 2021b; Wang et al., 2023) stands out. NGC utilizes the Causal Multilayer Perceptron (CMLP) model to train data, thereby enhancing the accuracy of causal inference tasks. By combining the predictive power of neural networks with the interpretability of Granger causality, NGC offers a robust framework for identifying causal relationships in complex datasets.
Janzing et al. (2012) introduced Information Geometric Causal Inference (IGCI) to address the nonlinear challenges encountered by the Additive Noise Model (ANM) algorithm. IGCI enhances causal inference by incorporating information entropy, providing a more effective method for dealing with nonlinear data structures. This approach allows for better differentiation between cause and effect in scenarios where traditional linear models fall short.
Mao et al. (2017) extended the application of the Convergent Cross Mapping (CCM) algorithm from bivariate to multivariate analysis by incorporating transfer entropy. This integration, named Phase State Delay Reconstruction - Transfer Entropy (PSDR-TE), effectively addresses the limitation of the CCM algorithm, which was originally designed for detecting bivariate relationships. PSDR-TE expands the capability of causal discovery to more complex multivariate systems, improving the detection of causal links across multiple variables.
Raghu et al. (2018) proposed the Scalable Causation Discovery Algorithm (SCDA), which combines structural equation modeling-based and conditional independence-based methods. SCDA provides a solution for mixed data containing both continuous and discrete sequences. By integrating these two methodologies, SCDA can handle a broader range of data types and improve the robustness of causal inference in heterogeneous datasets.
In summary, hybrid methods in causal discovery leverage the strengths of multiple algorithms to address their respective weaknesses. By combining techniques such as Granger causality, structural equation modeling, neural networks, and information entropy, these hybrid approaches offer powerful tools for uncovering causal relationships in diverse and complex datasets.
5 Empirical Study Design
The experimental framework is structured across four distinct phases. The inaugural phase involves conducting a comparative assessment of algorithms applied to synthesised datasets with specific features while concurrently evaluating a range of performance metrics. Subsequently, the second phase encompasses presenting and analysing outcomes derived from the initial stage to extract meaningful insights. Transitioning to the third phase, real-world datasets are engaged for testing, utilising the insights garnered in the preceding phase to ascertain the optimal algorithm. This stage aims to verify whether the test results are consistent with the predicted optimal algorithm. The fourth and final phase entails deploying diverse data processing and testing methodologies to ascertain the metadata of the time series datasets. This, in turn, facilitates the extrapolation of insights from the second phase to previously unexplored datasets.
5.1 Datasets
This experimental inquiry necessitates two dataset categories: synthetic datasets designed to explore underlying patterns and real-world datasets serving the purpose of validation. The data generation structure of the artificial dataset is shown in Figure 8.
Figure 8: Data generator structure
Given the categorization of testing algorithms into three classes, a concomitant preparation of diverse composite datasets becomes imperative. The first category entails a data generator tasked with producing causal pairs, whereby the data generation tool within the causal discovery toolbox (Kalainathan and Goudet, 2019b) was employed. The second and third categories encompass instantaneous and time-delay causality, for which the data generator in Tigramite framework, posited by Runge et al. (2023a), was harnessed for data synthesis. In the fourth category, characterized by i.i.d. data, we adopted the data generation model established within the gCastle package.
Subsequently, we considered the dataset sizes for both time series and i.i.d. data. For time series data, the experimental framework established distinct time series lengths ranging from 50 to 300 time points for small-scale datasets, 300 to 1000 for medium-scale datasets, and 1000 to 3000 for large-scale datasets. For i.i.d. data, the sizes were similarly categorized into small (50, 100, 150, 200, 250), medium (300, 440, 580, 720, 860), and large (1000, 1400, 1800, 2200, 2600).
The design of dataset attributes necessitates attention to causal relationships and noise distribution types. Causal relationships within this context are bifurcated into linear and nonlinear relationships. Linear relationships are generated through polynomial operations on dataset variables, while nonlinear relationships involve trigonometric operations. The noise distribution types encompass Gaussian noise with parameters defined by a mean of 0 and a standard deviation of 1, and uniform noise spanning the interval (0,1).
Building upon the framework above, five datasets are generated for each data size, each subjected ten times to mitigate runtime-induced biases, resulting in a total of 180 distinct datasets for comprehensive algorithm evaluation. All generated datasets adhere to the causal sufficiency assumption and possess stability, prerequisites fundamental to the operation of the algorithmic processes.
Regarding authentic datasets, scarcity in datasets featuring established ground truth, particularly within the domain of time series, is evident. Consequently, this study incorporates two verifiable datasets. The first is the “Tuebingen” dataset, which comprises 100 real cause-effect pairs. Additionally, the “fMRI” dataset, which aims to investigate the Blood Oxygen Level Dependent (BOLD) signal across 28 distinct intrinsic brain networks, is also integrated into the study.
5.2 Evaluation Metrics
In devising this project’s evaluation criteria, we tried to encompass multiple aspects as comprehensively as possible. Recognizing that a singular indicator might introduce bias, we select five distinct indicators for our assessment.
The initial metric utilized is the F1 score, which serves as a prevalent evaluation criterion in causal discovery due to its capacity to assess the model’s overall performance. The F1 score is computed as follows:
$F1=\frac{2\times precision\times recall}{precision+recall}$ (5)
To augment the evaluation of the model’s robustness, specifically its capacity to discern TPR and FPR, we employed the Area Under the Receiver Operating Characteristic (AUROC) as a metric, which can be obtained by computing the area under the ROC curve.
The first two indicators signify that higher numerical values correspond to better performance, with a preference to assess the accurate inference. For assessing the extent of false causality present in the model, we employ the False Positive Rate (FPR), computed as follows:
$FPR=\sum_{i}\frac{e_{i}}{E-E_{GT}},e_{i}\in E_{M}/E_{GT}$ (6)
Subsequently, we incorporate a graph-based causal discovery metric called Structural Hamming Distance (SHD). This metric directly reveals the number of incorrectly inferred edges by comparing the differences between the ground truth and the estimated causal graph.
Lastly, we meticulously recorded the run time of each algorithm, given that the time cost was considered a significant aspect of this experiment. This particular indicator will serve as a crucial criterion for assessing the efficiency of the algorithms.
5.3 Algorithms
By collecting resources on GitHub, we introduce in this section the algorithms selected for the experiment, as well as their source code and packages. We mainly use four packages, Causal Discovery Toolbox (CDT) (2019a), gCastle (2021b), causal_discovery_for_time_series (CD_TS)(2022), causal-learn (2024) to implement the testing algorithms.
As shown in Table 4, we select MVGC and PWGC, implemented in CD_TS(2022), from the repertoire of GC methods, as the subject of experimentation. It is worth noting that MVGC exclusively addresses delayed causality while PWGC addresses pairwise causality, so their applicability has certain limitations.
In the context of the conditional independence-based method, Runge et al. (2023b) developed PCMCI and its variant algorithms for time series. The oCSE, tsFCI and DYNOTEARS algorithms are also included in the experiment for time series. Considering i.i.d. data, we chooose PC, FCI, GES, GRaSP, ES and CDS. These methods are chosen for their robust performance in different scenarios, providing a comprehensive evaluation of algorithmic capabilities.
Most algorithms rooted in the state space dynamic-based approach primarily concentrate on resolving the directionality of causal pairs. Among these, we examine the classical CCM and PAI algorithms (Javier, 2021), categorizing them as instances of pairwise causality. Additionally, IGCI and ANM are incorporated as pairwise causality algorithms.
As for the structural equation modeling-based approach, our emphasis lies on exploring a variant of the LiNGAM algorithm (Ikeuchi et al., 2023) tailored for i.i.d. data and time series data, referred to as ICALiNGAM, DirectLiNGAM, RCD, CAM-UV and VARLiNGAM. The TiMINO algorithm is excluded from consideration due to its documented inferiority compared to PCMCI and TCDF (Nauta et al., 2019a).
TCDF algorithm(2019b) frequently emerges in diverse surveys and holds a pivotal position within the field, thus making it a suitable choice as a representative algorithm for deep learning-based methods. For i.i.d. data, we include three algorithms from gCastle package: DAG-GNN, CORL and GraNDAG. Besides, we do not include CGNN in this experiment due to its prolonged running time, which could impede the efficiency of the overall analysis.
Within the hybrid method, we opt to employ both NeuralGC (Tank et al., 2021a) and IGCI. NeuralGC possesses the ability to handle both delay and instantaneous causal relationships, while IGCI is specialized in resolving causal pairs.
In order to present a lucid exposition of the algorithms employed in this experiment, we marked the testing algorithms with * in Table 4. It is essential to highlight that, in pursuit of optimal performance for each algorithm, a series of individual tests were conducted. According to these tests, specific default values of several algorithms were modified to ensure the validation of the results.
Table 4: Causal Discovery Algorithms
Method Algorithm Time-series i.i.d. Faithfulness CMC Sufficiency Linear Software
Granger-based PWGC* CD_TS
MVGC* CD_TS
EGC
KGC KernelGrangerCausality
CopulaGC CopulaGrangerCausality
Condition independence-based oCSE* CD_TS
PC* CDT; gCastle; causal-learn
PCMCI* tigramite; CD_TS
FCI* causal-learn
tsFCI* CD_TS
GES* CDT; gCastle; causal-learn
GRaSP* causal-learn
ES* causal-learn
DYNOTEARS* CD_TS
CDS* CDT
State space dynamics-based CCM* causal_ccm
CMS
IOTA
PAI* causal_ccm
Structural equation modeling-based ICALiNGAM* LiNGAM; gCastle; causal-learn
DirectLiNGAM* LiNGAM; gCastle; causal-learn
VARLiNGAM* LiNGAM; CD_TS; causal-learn
RECI* CDT
RCD* LiNGAM; causal-learn
CAM-UV LiNGAM; causal-learn
ANM* CDT; gCastle; causal-learn
TiMINO* CD_TS
PNL gCastle; causal-learn
NOTEARS* gCastle
GOLEM* gCastle
Deep learning-based CGNN CDT
DAG-GNN* gCastle
TCDF* TCDF; CD_TS
ACD ACD
CORL* gCastle
GraNDAG* gCastle
Hybrid ARMA-LiNGAM LiNGAM
NeuralGC* Neural-GC
IGCI* CDT
SCDA
PSDR-TE
5.4 Environment Settings
Here we will expound upon the environment’s configuration of the entire code architecture, encompassing domains such as software provisioning, hardware parameters, database integration, and related facets.
The instantiation of this project is grounded in the Python programming language, with compilation facilitated through the PyCharm (professional edition) software. Upon successful compilation, the resultant code is subsequently uploaded onto the designated server for operational deployment.
The particular details about the server infrastructure are outlined herewith: The server infrastructure is established on the Ubuntu operating system, boasting four Graphics Processing Units (GPUs) that operate in tandem. These GPUs are compatible with the 11.4 version of the Compute Unified Device Architecture (CUDA). Notably, each GPU has a computational prowess of 350 Watts and 24,268 Megabytes of memory.
The source code’s comprehensive architecture encompasses five submodules, each fulfilling designated functions. The initial submodule, termed the “dataset”, functionality encompasses the storage of datasets and ground truth sets in CSV format alongside the capacity to generate synthetic datasets. Moving forward, the second submodule, designated the “examples”, serves the dual purpose of harmonizing testing algorithms and facilitating the visual representation of obtained outcomes. The third submodule, “save”, is dedicated to the archival of causal discovery results derived from algorithmic testing endeavours. Within the fourth submodule, designated the “src”, the central objective pertains to integrating external configuration files, thus ensuring seamless operational functionality of the codebase. The final submodule, aptly labelled the “doc”, assumes the role of elucidating practical usage instances of this code library while presenting information regarding the version of the installation package.
5.5 Research Questions
With a focus on the analysis of algorithm performance and result effectiveness, our experiment addresses the following research questions:
RQ 1 (Comparison of Algorithm Performance): Among the assessed algorithms, which one demonstrates superior effectiveness or efficiency under specific data characteristics?
RQ 1 is the baseline question in the experiment and a prerequisite for answering other questions. By evaluating the impact of data features on algorithm performance, we establish a foundation that informs subsequent steps in our experimental analysis.
RQ 2 (Real-World Applicability): Are the insights derived from the synthetic datasets consistent with those acquired from the real datasets?
RQ 2 is crucial for determining the effectiveness of insights gained from RQ 1, as it connects the findings from synthetic data to real-world scenarios. This question ensures that our conclusions are not limited to controlled experimental conditions but are also valid in practical applications.
RQ 3 (Generalization to Unknown Datasets):
  • RQ 3.1 (Metadata Recognition for Algorithm Selection): Is it feasible to precisely capture the representative attributes of unknown datasets using their metadata to ascertain the optimal algorithm based on our previous conclusions?
  • RQ 3.2 (Practical Recommendations for Users): Upon successfully identifying the optimal algorithm for an unknown dataset in RQ 3.1, what practical recommendations can we provide to users for selecting appropriate methods for their specific datasets?
RQ 3 is a challenging aspect of the experiment as it involves extending the results of RQ 1 and RQ 2 to a broader range of applications. RQ 3.1 focuses on the feasibility of applying our findings to new datasets by analyzing their metadata, ensuring that our methods are robust and versatile. RQ 3.2 aims to translate these validated approaches into practical, user-friendly guidelines that assist practitioners in choosing the best algorithms for their unique datasets, thus bridging the gap between theoretical research and practical implementation.
6 Results Analysis
Within this section, the research questions posited in the preceding section will be addressed systematically, resulting in generalizable findings according to experimental plots and tables.
6.1 Answer to RQ1: Comparison of Algorithm Performance
Through meticulous experimentation, we have derived comparative graphs and ranking tables for algorithms across four distinct causality types. Considering that the ranking algorithm table only displays the average of metrics across all data sizes, we still need violin plots in Appendix A to supplement the changes in recommendation algorithms under specific sample lengths. This section will subsequently present an in-depth analysis of each category.
The causal relationships, whether linear or nonlinear, in conjunction with noise distributions (Gaussian or non-Gaussian), yield four distinct subgraphs, each representing different data modalities. Dataset time lengths are categorized as follows: small (50, 300), medium (300, 1000), and large (1000, 5000). Each dataset is processed ten times to compute mean values.
6.1.1 Pairwise causality
Figure A illustrates the experimental outcomes for the pairwise causal discovery algorithms. We identify an algorithm that consistently performs well as the reference algorithm. For instance, within linear relationships, the CCM algorithm serves as the benchmark, whereas the PAI algorithm is selected for nonlinear relationships. The deviations between the algorithms and their respective reference counterparts are depicted using violin plots in Appendix A.
Acknowledging the limitations of violin plots in conveying precise numerical results, we supplement our analysis with a ranking table. Table 5 quantifies and compares the performance disparities among the algorithms under consideration. The algorithms are ordered in descending sequence based on metric values in this table. Higher F1 scores and AUROC values indicate superior model performance, while lower FPR and SHD values reflect better algorithmic behavior.
Table 5: Ranking table of pairwise causality
Linear Gaussian Linear Non-Gaussian Non-Linear Gaussian Non-Linear Non-Gaussian
F1$\uparrow$ CCM $(0.56\pm 0.04)$ CCM $(0.57\pm 0.06)$ PAI $(0.71\pm 0.05)$ PAI $(0.73\pm 0.07)$
PAI $(0.51\pm 0.07)$ PAI $(0.56\pm 0.09)$ CDS $(0.59\pm 0.16)$ CCM $(0.65\pm 0.08)$
PWGC $(0.50\pm 0.03)$ PWGC $(0.53\pm 0.03)$ CCM $(0.57\pm 0.04)$ CDS $(0.59\pm 0.16)$
CDS $(0.47\pm 0.13)$ CDS $(0.50\pm 0.14)$ PWGC $(0.48\pm 0.04)$ PWGC $(0.50\pm 0.03)$
IGCI $(0.44\pm 0.04)$ IGCI $(0.48\pm 0.04)$ ANM $(0.36\pm 0.07)$ ANM $(0.35\pm 0.07)$
ANM $(0.33\pm 0.09)$ ANM $(0.31\pm 0.09)$ IGCI $(0.32\pm 0.05)$ IGCI $(0.31\pm 0.05)$
RECI $(0.29\pm 0.03)$ RECI $(0.21\pm 0.04)$ RECI $(0.22\pm 0.03)$ RECI $(0.21\pm 0.03)$
[…]
PAI $(0.63\pm 0.05)$ PAI $(0.67\pm 0.07)$ CDS $(0.72\pm 0.04)$ CCM $(0.73\pm 0.06)$
CDS $(0.62\pm 0.03)$ CDS $(0.65\pm 0.03)$ CCM $(0.68\pm 0.03)$ CDS $(0.71\pm 0.06)$
PWGC $(0.62\pm 0.03)$ PWGC $(0.64\pm 0.05)$ PWGC $(0.61\pm 0.06)$ PWGC $(0.62\pm 0.03)$
ANM $(0.59\pm 0.04)$ IGCI $(0.61\pm 0.03)$ ANM $(0.59\pm 0.04)$ ANM $(0.59\pm 0.04)$
IGCI $(0.58\pm 0.03)$ ANM $(0.59\pm 0.03)$ IGCI $(0.49\pm 0.04)$ IGCI $(0.49\pm 0.04)$
RECI $(0.47\pm 0.02)$ RECI $(0.45\pm 0.03)$ RECI $(0.42\pm 0.02)$ RECI $(0.41\pm 0.02)$
[…]
IGCI $(0.56\pm 0.04)$ IGCI $(0.52\pm 0.04)$ IGCI $(0.69\pm 0.05)$ IGCI $(0.69\pm 0.05)$
PWGC $(0.50\pm 0.03)$ PWGC $(0.47\pm 0.03)$ PWGC $(0.52\pm 0.04)$ PWGC $(0.50\pm 0.03)$
PAI $(0.49\pm 0.07)$ PAI $(0.45\pm 0.09)$ CCM $(0.43\pm 0.04)$ CCM $(0.35\pm 0.08)$
CDS $(0.47\pm 0.13)$ CCM $(0.43\pm 0.06)$ ANM $(0.37\pm 0.04)$ CDS $(0.34\pm 0.10)$
CCM $(0.44\pm 0.04)$ CDS $(0.42\pm 0.11)$ CDS $(0.35\pm 0.11)$ ANM $(0.34\pm 0.07)$
ANM $(0.31\pm 0.05)$ ANM $(0.28\pm 0.08)$ PAI $(0.29\pm 0.05)$ PAI $(0.27\pm 0.07)$
[…]
ANM $(6.68\pm 0.86)$ ANM $(6.87\pm 0.89)$ IGCI $(6.85\pm 0.48)$ IGCI $(6.86\pm 0.56)$
IGCI $(5.57\pm 0.35)$ IGCI $(5.24\pm 0.36)$ ANM $(6.42\pm 0.72)$ ANM $(6.54\pm 0.73)$
CDS $(5.32\pm 1.28)$ CDS $(4.96\pm 1.40)$ PWGC $(5.22\pm 0.41)$ PWGC $(4.97\pm 0.33)$
PWGC $(4.97\pm 0.32)$ PWGC $(4.66\pm 0.34)$ CCM $(4.27\pm 0.42)$ CDS $(4.13\pm 1.60)$
PAI $(4.93\pm 0.72)$ PAI $(4.45\pm 0.93)$ CDS $(4.15\pm 1.64)$ CCM $(3.54\pm 0.75)$
CCM $(4.39\pm 0.39)$ CCM $(4.34\pm 0.62)$ PAI $(2.88\pm 0.54)$ PAI $(2.71\pm 0.68)$
Drawing insights from Figure A and Table 5, we synthesize a comprehensive summary delineating the preeminent algorithmic selections under diverse application scenarios. In scenarios where the overall performance of the model is emphasized, the F1 score assumes paramount importance. For the linear dataset, CCM is advocated as the optimal choice, while the RECI algorithm, deemed the least effective, exhibits an F1 score approximately 50% lower than that of CCM. Across other datasets, the PAI algorithm consistently emerges as the pinnacle performer. Particularly in nonlinear and Gaussian datasets, PAI outperforms its peers significantly, with its F1 score surpassing that of the second-best algorithm by 20%. Note that the optimal algorithm for each data size is always among the top three algorithms in the average ranking, which means that the sample length has little impact on our recommendation for F1.
In the pursuit of heightened system robustness, AUROC is prioritized. The recommendations from this scenario align with those derived from the F1 score, reaffirming the advisability of employing either the CCM or PAI algorithms. However, the PAI algorithm is not competitive in small sample sizes of linear datasets and is only recommended in medium or large sample sizes ($L>300$).
For systems sensitive to false causality, preference should be given to the FPR metric. In cases involving linear interrelationships among variables, ANM is the preferred choice, with the worst algorithm’s FPR being more than twice as high as that of ANM. CDS is more recommended in small size datasets. Conversely, when dealing with nonlinear relationships among variables, the PAI algorithm emerges as the optimal selection.
When dealing with situations characterized by limited error tolerance, SHD takes precedence. The distinction between the CCM and PAI algorithms is minimal within linear datasets. However, in nonlinear datasets, PAI demonstrates a noteworthy reduction in SHD, exceeding a minimum of 20% compared to other algorithms.
6.1.2 Instantaneous causality
Following this exposition, we elucidate the performance evaluation of instantaneous causal discovery algorithms, as visually depicted in Figure A and quantitatively delineated in Table 6. Figure A records the performance of seven algorithms, with oCSE serving as the reference algorithm for Gaussian noise datasets and VARLiNGAM as the reference for non-Gaussian noise datasets.
Similar to the analysis of pairwise causality, we consider five scenarios when evaluating instantaneous causality algorithms. The top 3 best-performing algorithms under each metric can be summarized from Table 6. We will present them in Table 10 and not elaborate here. It is necessary to supplement some insights based on data sample size in Figure A.
Firstly, when prioritizing F1 scores, for non-linear Gaussian datasets, tsFCI performs in the top three in small data sizes, while PCMCI performs better in large data sizes. For Nonlinear non-Gaussian datasets, VARLiNGAM will only be recommended when the sample length is small ($L<1000$). The average ranking of algorithm performance under AUROC metric is basically consistent with the ranking under each data size.
In the context of prioritizing FPR metrics, PCMCI belongs to the recommendation algorithms in all four data types, indicating its superior performance under the FPR metric. For Linear Gaussian datasets, tsFCI is only recommended when the sample size is small.
Under the SHD metric, oCSE, PCMCI, and VARLiNGAM are recommended for all four data types, proving that these three algorithms have stable performance to find the real causal relations and are not heavily dependent on data features.
Table 6: Ranking table of instantaneous causality
Linear Gaussian Linear Non-Gaussian Non-Linear Gaussian Non-Linear Non-Gaussian
F1$\uparrow$ oCSE$(0.74\pm 0.23)$ VARLiNGAM$(0.88\pm 0.13)$ NeuralGC $(0.53\pm 0.10)$ NeuralGC $(0.52\pm 0.07)$
PCMCI $(0.71\pm 0.20)$ oCSE $(0.72\pm 0.18)$ tsFCI $(0.45\pm 0.08)$ VARLiNGAM $(0.50\pm 0.15)$
VARLiNGAM $(0.63\pm 0.07)$ PCMCI $(0.71\pm 0.16)$ PCMCI $(0.44\pm 0.20)$ tsFCI $(0.48\pm 0.10)$
NeuralGC $(0.61\pm 0.12)$ DYNOTEARS $(0.64\pm 0.10)$ oCSE $(0.42\pm 0.21)$ PCMCI $(0.41\pm 0.12)$
DYNOTEARS $(0.56\pm 0.12)$ NeuralGC $(0.61\pm 0.12)$ VARLiNGAM $(0.36\pm 0.15)$ oCSE $(0.40\pm 0.14)$
tsFCI $(0.50\pm 0.07)$ tsFCI $(0.50\pm 0.05)$ DYNOTEARS $(0.31\pm 0.15)$ DYNOTEARS $(0.37\pm 0.13)$
TCDF $(0.32\pm 0.16)$ TCDF $(0.39\pm 0.12)$ TCDF $(0.22\pm 0.15)$ TCDF $(0.24\pm 0.11)$
[…]
PCMCI $(0.77\pm 0.14)$ oCSE $(0.79\pm 0.12)$ oCSE $(0.59\pm 0.13)$ tsFCI $(0.58\pm 0.10)$
NeuralGC $(0.72\pm 0.07)$ PCMCI $(0.78\pm 0.11)$ PCMCI $(0.58\pm 0.14)$ oCSE $(0.58\pm 0.08)$
VARLiNGAM $(0.70\pm 0.06)$ NeuralGC $(0.71\pm 0.07)$ tsFCI $(0.56\pm 0.08)$ VARLiNGAM $(0.57\pm 0.14)$
tsFCI $(0.64\pm 0.05)$ DYNOTEARS $(0.70\pm 0.10)$ VARLiNGAM $(0.45\pm 0.14)$ PCMCI $(0.41\pm 0.12)$
DYNOTEARS $(0.63\pm 0.10)$ tsFCI $(0.63\pm 0.04)$ DYNOTEARS $(0.41\pm 0.15)$ DYNOTEARS $(0.47\pm 0.14)$
TCDF $(0.41\pm 0.14)$ TCDF $(0.48\pm 0.09)$ TCDF $(0.31\pm 0.13)$ TCDF $(0.34\pm 0.10)$
[…]
TCDF $(0.70\pm 0.36)$ tsFCI $(0.60\pm 0.33)$ TCDF $(0.81\pm 0.28)$ TCDF $(0.69\pm 0.22)$
DYNOTEARS $(0.63\pm 0.42)$ TCDF $(0.40\pm 0.23)$ VARLiNGAM $(0.76\pm 0.35)$ DYNOTEARS $(0.66\pm 0.39)$
VARLiNGAM $(0.62\pm 0.25)$ DYNOTEARS $(0.37\pm 0.32)$ DYNOTEARS $(0.75\pm 0.34)$ VARLiNGAM$(0.54\pm 0.41)$
tsFCI $(0.61\pm 0.35)$ PCMCI $(0.23\pm 0.14)$ tsFCI $(0.57\pm 0.28)$ tsFCI $(0.48\pm 0.30)$
PCMCI $(0.24\pm 0.17)$ oCSE $(0.22\pm 0.16)$ PCMCI $(029\pm 0.15)$ PCMCI $(0.29\pm 0.17)$
oCSE $(0.22\pm 0.19)$ VARLiNGAM $(0.21\pm 0.29)$ oCSE $(0.25\pm 0.16)$ oCSE $(0.22\pm 0.19)$
[…]
NeuralGC $(6.37\pm 2.60)$ NeuralGC $(6.19\pm 2.27)$ NeuralGC $(7.20\pm 2.00)$ NeuralGC $(7.25\pm 1.96)$
TCDF $(6.02\pm 0.98)$ TCDF $(5.76\pm 0.74)$ DYNOTEARS $(6.80\pm 1.44)$ TCDF $(6.56\pm 0.77)$
DYNOTEARS $(5.67\pm 1.59)$ DYNOTEARS $(4.45\pm 1.59)$ TCDF $(6.77\pm 1.29)$ DYNOTEARS $(5.97\pm 1.62)$
VARLiNGAM $(5.24\pm 0.99)$ PCMCI $(2.88\pm 1.18)$ VARLiNGAM $(6.44\pm 1.62)$ PCMCI $(5.37\pm 0.71)$
PCMCI $(2.92\pm 1.53)$ oCSE $(2.73\pm 1.39)$ oCSE $(4.91\pm 1.03)$ VARLiNGAM $(5.33\pm 1.71)$
oCSE $(2.57\pm 1.60)$ VARLiNGAM $(1.81\pm 1.98)$ PCMCI $(4.91\pm 1.10)$ oCSE $(5.21\pm 0.69)$
6.1.3 Time-delay Causality
Below, we will analyze the performance comparison of time-delay causal algorithms in detail. The top 3 best-performing algorithms under each metric can be summarized from Table 7. We will present them in Table 10 and not elaborate here.
Table 7: Ranking table of time-delay causality
Linear Gaussian Linear Non-Gaussian Non-Linear Gaussian Non-Linear Non-Gaussian
F1$\uparrow$ VARLiNGAM $(0.87\pm 0.14)$ PCMCI$(0.88\pm 0.11)$ PCMCI $(0.78\pm 0.08)$ PCMCI $(0.79\pm 0.07)$
MVGC $(0.86\pm 0.09)$ VARLiNGAM $(0.87\pm 0.13)$ DYNOTEARS $(0.75\pm 0.07)$ NeuralGC $(0.77\pm 0.09)$
DYNOTEARS $(0.86\pm 0.11)$ DYNOTEARS $(0.87\pm 0.11)$ VARLiNGAM $(0.74\pm 0.07)$ DYNOTEARS $(0.74\pm 0.09)$
PCMCI $(0.86\pm 0.11)$ MVGC $(0.87\pm 0.11)$ NeuralGC $(0.74\pm 0.07)$ oCSE $(0.73\pm 0.09)$
NeuralGC $(0.83\pm 0.13)$ NeuralGC $(0.83\pm 0.14)$ MVGC $(0.72\pm 0.06)$ VARLiNGAM $(0.73\pm 0.07)$
oCSE $(0.78\pm 0.15)$ oCSE $(0.81\pm 0.12)$ oCSE $(0.71\pm 0.10)$ MVGC $(0.72\pm 0.08)$
TCDF $(0.71\pm 0.12)$ TCDF $(0.73\pm 0.10)$ tsFCI $(0.67\pm 0.06)$ tsFCI $(0.67\pm 0.04)$
tsFCI $(0.68\pm 0.07)$ tsFCI $(0.69\pm 0.05)$ TCDF $(0.60\pm 0.07)$ TCDF $(0.60\pm 0.10)$
[…]
VARLiNGAM $(0.92\pm 0.08)$ VARLiNGAM $(0.93\pm 0.08)$ MVGC $(0.82\pm 0.05)$ MVGC $(0.84\pm 0.05)$
DYNOTEARS $(0.90\pm 0.09)$ PCMCI $(0.91\pm 0.07)$ DYNOTEARS $(0.81\pm 0.07)$ NeuralGC $(0.82\pm 0.07)$
PCMCI $(0.90\pm 0.07)$ DYNOTEARS $(0.91\pm 0.09)$ VARLiNGAM $(0.81\pm 0.05)$ oCSE $(0.82\pm 0.06)$
NeuralGC $(0.87\pm 0.10)$ oCSE $(0.87\pm 0.07)$ NeuralGC $(0.80\pm 0.06)$ DYNOTEARS $(0.81\pm 0.07)$
oCSE $(0.86\pm 0.08)$ NeuralGC $(0.87\pm 0.10)$ oCSE $(0.79\pm 0.08)$ VARLiNGAM $(0.80\pm 0.05)$
TCDF $(0.82\pm 0.09)$ TCDF $(0.84\pm 0.08)$ tsFCI $(0.74\pm 0.04)$ tsFCI $(0.75\pm 0.02)$
tsFCI $(0.76\pm 0.03)$ tsFCI $(0.77\pm 0.03)$ TCDF $(0.73\pm 0.07)$ TCDF $(0.72\pm 0.09)$
[…]
NeuralGC $(0.45\pm 0.39)$ NeuralGC $(0.43\pm 0.41)$ TCDF $(0.50\pm 0.21)$ TCDF $(0.54\pm 0.26)$
oCSE $(0.33\pm 0.32)$ oCSE $(0.31\pm 0.30)$ tsFCI $(0.46\pm 0.20)$ DYNOTEARS $(0.51\pm 0.35)$
DYNOTEARS $(0.26\pm 0.29)$ DYNOTEARS $(0.25\pm 0.35)$ DYNOTEARS $(0.46\pm 0.35)$ tsFCI $(0.49\pm 0.23)$
PCMCI $(0.25\pm 0.19)$ PCMCI $(0.21\pm 0.17)$ PCMCI $(0.45\pm 0.16)$ PCMCI$(0.47\pm 0.21)$
TCDF $(0.24\pm 0.25)$ TCDF $(0.18\pm 0.20)$ oCSE $(0.41\pm 0.23)$ VARLiNGAM $(0.45\pm 0.21)$
VARLiNGAM $(0.11\pm 0.12)$ VARLiNGAM $(0.09\pm 0.13)$ VARLiNGAM $(0.36\pm 0.24)$ oCSE $(0.37\pm 0.23)$
MVGC $(0.07\pm 0.24)$ MVGC $(0.04\pm 0.13)$ MVGC $(0.25\pm 0.25)$ MVGC $(0.22\pm 0.22)$
[…]
TCDF $(4.58\pm 1.68)$ TCDF $(4.35\pm 1.33)$ TCDF $(6.10\pm 0.92)$ TCDF $(6.39\pm 1.33)$
oCSE $(4.15\pm 2.22)$ NeuralGC $(4.12\pm 3.59)$ NeuralGC $(5.86\pm 1.90)$ NeuralGC $(5.38\pm 2.41)$
NeuralGC $(4.14\pm 3.33)$ oCSE $(3.62\pm 1.74)$ oCSE $(5.54\pm 1.52)$ VARLiNGAM $(5.34\pm 1.14)$
DYNOTEARS $(2.91\pm 2.47)$ DYNOTEARS $(2.68\pm 2.60)$ VARLiNGAM $(5.31\pm 1.17)$ DYNOTEARS $(5.25\pm 2.02)$
PCMCI $(2.80\pm 1.95)$ MVGC $(2.43\pm 1.97)$ MVGC $(5.25\pm 0.85)$ oCSE $(5.06\pm 1.58)$
MVGC $(2.49\pm 1.60)$ PCMCI $(2.30\pm 1.80)$ DYNOTEARS $(5.10\pm 1.62)$ MVGC $(4.97\pm 1.09)$
VARLiNGAM $(2.29\pm 1.98)$ VARLiNGAM $(2.18\pm 2.06)$ PCMCI $(4.41\pm 1.00)$ PCMCI $(4.20\pm 1.11)$
It is necessary to supplement some insights based on data sample size in Figure A. Firstly, we analyze the algorithms under the F1 metric. For Linear Gaussian data, VARLiNGAM performs better in larger sample sizes. On the contrary, MVGC is more suitable for small sample size data. The top three ranking algorithms perform relatively stably on Linear non-Gaussian and Nonlinear Gaussian data, and their rankings do not change significantly based on changes in sample size. For Nonlinear non-Gaussian data, DYNOTEARS is recommended only when the sample size is small, since it does not ranked in the top three in large sample size.
The ordering of algorithms according to the AUROC metric exhibits a relatively stable pattern. Specifically, the top three algorithms in average ranking table reflects the recommended algorithms for AUROC across all data sizes.
In evaluating performance utilizing the FPR metric, for linear data, TCDF ranks in the top three only when the sample length $>1000$, so it is not a recommended algorithm when the sample size is small. For non-linear data, VARLiNGAM is not outstanding when the sample length $<300$ and is only recommended for use in large data size. Note that although the oCSE algorithm does not always in the top three of the mean ranking, it is always the best algorithm in small sample sizes ($L<300$).
Considering SHD, for linear type, VARLiNGAM is more suitable for large sample data, while PCMCI is more suitable for small sample data. For nonlinear type, MVGC is more suitable for small sample data. In Nonlinear non-Gaussian data, although NeuralGC is not among the top three algorithms, it is the best algorithm when sample length $>1000$ and can be recommended as a supplementary algorithm.
After comparing the performance of time-delay causal discovery algorithms, it can be concluded that tsFCI are not recommended in any scenario, as they are not competitive across any metric.
6.1.4 i.i.d. Causality
Below, we will analyze the performance comparison of i.i.d. data causal algorithms in detail.
Table 8: Ranking table of i.i.d. causality
Linear Gaussian Linear Non-Gaussian Non-Linear Gaussian Non-linear Non-Gaussian
F1$\uparrow$ GOLEM $(0.99\pm 0.02)$ DirectLiNGAM$(0.99\pm 0.05)$ GraNDAG $(0.51\pm 0.24)$ CORL $(0.51\pm 0.14)$
CORL $(0.98\pm 0.04)$ CORL $(0.98\pm 0.04)$ CORL $(0.49\pm 0.07)$ GOLEM $(0.46\pm 0.11)$
DAG-GNN $(0.95\pm 0.08)$ GOLEM $(0.95\pm 0.07)$ GES $(0.44\pm 0.11)$ GraNDAG $(0.39\pm 0.18)$
NOTEARS $(0.90\pm 0.10)$ ICALiNGAM $(0.94\pm 0.09)$ GOLEM $(0.44\pm 0.05)$ ICALiNGAM $(0.36\pm 0.15)$
ES $(0.85\pm 0.11)$ DAG-GNN $(0.91\pm 0.09)$ RCD $(0.43\pm 0.06)$ DAG-GNN $(0.35\pm 0.12)$
ICALiNGAM $(0.70\pm 0.13)$ RCD $(0.87\pm 0.17)$ FCI $(0.41\pm 0.12)$ RCD $(0.33\pm 0.06)$
GRaSP $(0.62\pm 0.13)$ NOTEARS $(0.86\pm 0.09)$ PC $(0.40\pm 0.12)$ GES $(0.30\pm 0.06)$
PC $(0.57\pm 0.10)$ GRaSP $(0.66\pm 0.15)$ ICALiNGAM $(0.38\pm 0.08)$ PC $(0.27\pm 0.09)$
FCI $(0.51\pm 0.16)$ PC $(0.58\pm 0.10)$ ES $(0.35\pm 0.12)$ NOTEARS $(0.25\pm 0.11)$
GES $(0.46\pm 0.14)$ FCI $(0.51\pm 0.13)$ NOTEARS $(0.32\pm 0.08)$ FCI $(0.24\pm 0.08)$
DirectLiNGAM $(0.40\pm 0.09)$ ES $(0.51\pm 0.13)$ DAG-GNN $(0.31\pm 0.06)$ ES $(0.23\pm 0.08)$
GraNDAG $(0.36\pm 0.24)$ GES $(0.49\pm 0.13)$ GRaSP $(0.29\pm 0.18)$ GRaSP $(0.15\pm 0.07)$
RCD $(0.35\pm 0.04)$ GraNDAG $(0.18\pm 0.16)$ DirectLiNGAM $(0.11\pm 0.06)$ DirectLiNGAM $(0.12\pm 0.08)$
[…]
CORL $(0.98\pm 0.03)$ CORL $(0.98\pm 0.04)$ NOTEARS $(0.63\pm 0.04)$ GOLEM $(0.58\pm 0.09)$
DAG-GNN $(0.96\pm 0.07)$ GOLEM $(0.96\pm 0.06)$ GraNDAG $(0.62\pm 0.23)$ DAG-GNN $(0.53\pm 0.09)$
NOTEARS $(0.91\pm 0.09)$ ICALiNGAM $(0.95\pm 0.08)$ ICALiNGAM $(0.60\pm 0.05)$ NOTEARS $(0.52\pm 0.08)$
ES $(0.86\pm 0.09)$ DAG-GNN $(0.92\pm 0.08)$ DAG-GNN $(0.59\pm 0.05)$ ICALiNGAM $(0.48\pm 0.12)$
ICALiNGAM $(0.73\pm 0.12)$ RCD $(0.89\pm 0.14)$ GOLEM $(0.56\pm 0.07)$ GraNDAG $(0.47\pm 0.19)$
GRaSP $(0.67\pm 0.13)$ NOTEARS $(0.88\pm 0.09)$ RCD $(0.52\pm 0.07)$ RCD $(0.46\pm 0.06)$
FCI $(0.61\pm 0.12)$ GRaSP $(0.71\pm 0.13)$ GES $(0.49\pm 0.10)$ GES $(0.35\pm 0.06)$
PC $(0.61\pm 0.09)$ FCI $(0.63\pm 0.08)$ FCI $(0.49\pm 0.09)$ GRaSP $(0.33\pm 0.10)$
RCD $(0.60\pm 0.02)$ PC $(0.61\pm 0.09)$ GRaSP $(0.48\pm 0.11)$ PC $(0.31\pm 0.08)$
GES $(0.52\pm 0.13)$ ES $(0.56\pm 0.12)$ PC $(0.46\pm 0.11)$ FCI $(0.30\pm 0.08)$
GraNDAG $(0.47\pm 0.23)$ GES $(0.54\pm 0.12)$ ES $(0.40\pm 0.11)$ ES $(0.28\pm 0.08)$
DirectLiNGAM $(0.45\pm 0.08)$ GraNDAG $(0.29\pm 0.19)$ DirectLiNGAM $(0.27\pm 0.08)$ DirectLiNGAM $(0.25\pm 0.12)$
[…]
GES $(0.50\pm 0.19)$ GES $(0.42\pm 0.17)$ GraNDAG $(0.24\pm 0.34)$ GraNDAG $(0.28\pm 0.34)$
DirectLiNGAM $(0.41\pm 0.10)$ PC $(0.18\pm 0.06)$ PC $(0.22\pm 0.06)$ PC $(0.22\pm 0.13)$
ICALiNGAM $(0.23\pm 0.11)$ GRaSP $(0.12\pm 0.09)$ GES $(0.21\pm 0.06)$ GES$(0.22\pm 0.15)$
PC $(0.17\pm 0.05)$ GraNDAG $(0.10\pm 0.06)$ RCD $(0.19\pm 0.14)$ RCD $(0.21\pm 0.16)$
GRaSP $(0.16\pm 0.12)$ RCD $(0.09\pm 0.13)$ FCI $(0.12\pm 0.05)$ FCI $(0.16\pm 0.09)$
GraNDAG $(0.12\pm 0.11)$ DAG-GNN $(0.06\pm 0.07)$ GOLEM $(0.11\pm 0.07)$ GRaSP $(0.09\pm 0.09)$
ES $(0.09\pm 0.06)$ FCI $(0.05\pm 0.04)$ GRaSP $(0.09\pm 0.06)$ GOLEM $(0.06\pm 0.05)$
FCI $(0.06\pm 0.04)$ NOTEARS $(0.05\pm 0.05)$ DirectLiNGAM $(0.07\pm 0.03)$ DirectLiNGAM $(0.05\pm 0.02)$
DAG-GNN $(0.05\pm 0.09)$ GOLEM $(0.04\pm 0.07)$ CORL $(0.04\pm 0.05)$ CORL $(0.05\pm 0.04)$
NOTEARS $(0.04\pm 0.05)$ ICALiNGAM $(0.04\pm 0.05)$ ICALiNGAM $(0.04\pm 0.02)$ ICALiNGAM $(0.05\pm 0.03)$
CORL $(0.02\pm 0.05)$ CORL $(0.02\pm 0.06)$ DAG-GNN $(0.03\pm 0.03)$ DAG-GNN $(0.04\pm 0.03)$
GOLEM $(0.01\pm 0.02)$ DirectLiNGAM $(0.01\pm 0.03)$ NOTEARS $(0.01\pm 0.01)$ NOTEARS $(0.02\pm 0.02)$
[…]
GES $(18.72\pm 5.41)$ GraNDAG $(15.75\pm 2.16)$ ES $(17.13\pm 3.39)$ GraNDAG $(13.95\pm 10.57)$
DirectLiNGAM $(17.45\pm 3.51)$ ES $(15.35\pm 3.93)$ DirectLiNGAM $(15.20\pm 0.60)$ ES $(13.93\pm 4.61)$
GraNDAG $(14.20\pm 4.60)$ FCI $(11.03\pm 1.75)$ FCI $(14.81\pm 1.11)$ GES $(12.97\pm 4.03)$
FCI $(10.96\pm 2.02)$ PC $(10.84\pm 2.21)$ GraNDAG $(13.91\pm 9.08)$ PC $(12.63\pm 3.56)$
PC $(10.76\pm 2.05)$ GRaSP $(8.48\pm 3.33)$ GES $(13.55\pm 2.26)$ FCI $(11.79\pm 3.41)$
GRaSP $(9.83\pm 3.35)$ RCD $(4.43\pm 6.60)$ GRaSP $(13.33\pm 2.11)$ GRaSP $(9.65\pm 2.21)$
ICALiNGAM $(8.27\pm 4.09)$ NOTEARS $(3.35\pm 2.54)$ PC $(13.24\pm 2.22)$ DirectLiNGAM $(9.13\pm 1.91)$
ES $(3.40\pm 2.35)$ DAG-GNN $(2.72\pm 0.2.77)$ DAG-GNN $(13.21\pm 0.75)$ NOTEARS $(8.31\pm 1.67)$
NOTEARS $(2.56\pm 2.70)$ GOLEM $(1.63\pm 2.16)$ GOLEM $(13.19\pm 1.76)$ DAG-GNN $(7.95\pm 1.73)$
DAG-GNN $(1.56\pm 2.79)$ ICALiNGAM $(1.43\pm 2.25)$ NOTEARS $(12.93\pm 0.90)$ ICALiNGAM $(7.77\pm 1.91)$
CORL $(0.56\pm 1.39)$ CORL $(0.67\pm 1.61)$ ICALiNGAM $(12.24\pm 1.05)$ GOLEM $(7.64\pm 1.85)$
GOLEM $(0.33\pm 0.71)$ DirectLiNGAM $(0.36\pm 1.35)$ CORL $(10.95\pm 1.39)$ CORL $(6.84\pm 1.99)$
Based on Figure A and Table 8, we can derive insights into the performance of recommendation algorithms across different i.i.d. data types. Analyzing the algorithms using the F1 metric, Table 8 reveals that the optimal algorithms vary according to specific data characteristics, with the CORL algorithm consistently ranking among the top three performers.
When considering the AUROC metric, the CORL algorithm shows superior performance on nonlinear datasets, although the improvement over the second-best algorithm is less than 10%. In contrast, GOLEM or DirectLiNGAM exhibit a slight advantage over CORL when applied to linear datasets.
Assessing performance using the FPR metric, NOTEARS emerges as the most proficient algorithm for nonlinear datasets. The GraNDAG algorithm performs best when the sample size is large, so we also include it in Table 8. For linear datasets, GOLEM performs best under Gaussian noise distribution, while DirectLiNGAM excels with non-Gaussian noise. Note that we need to supplement the DAG-GNN algorithm on large datasets, although it performs poorly on small-sized datasets.
Evaluated using the SHD metric, CORL demonstrates optimal performance on nonlinear datasets. Conversely, GOLEM achieves the lowest SHD values on linear datasets with Gaussian noise, while DirectLiNGAM performs best with non-Gaussian noise.
6.1.5 Discussion on Algorithm Efficiency
It is known that effectiveness does not equal efficiency. Even if some algorithms have good causal discovery performance, they may not be suitable for users because of long running time. So here we will specifically discuss the running time of the algorithms we tested in the first four sections, as shown in Table 9, to help users make a trade-off between effectiveness and efficiency.
Table 9: Ranking table of run time (s)$\downarrow$
Linear Gaussian Linear Non-Gaussian Non-Linear Gaussian Non-Linear Non-Gaussian
Pairwise ANM $(13.19\pm 24.70)$ ANM $(13.22\pm 24.77)$ ANM $(13.34\pm 25.09)$ ANM $(13.20\pm 24.85)$
CCM $(6.03\pm 10.49)$ CCM $(6.03\pm 10.50)$ CCM $(5.99\pm 10.40)$ CCM $(6.00\pm 10.45)$
PAI $(6.02\pm 10.48)$ PAI $(6.02\pm 10.48)$ PAI $(5.98\pm 10.40)$ PAI $(5.98\pm 10.40)$
CDS $(0.68\pm 0.39)$ CDS $(0.68\pm 0.39)$ CDS $(0.75\pm 0.48)$ CDS $(0.75\pm 0.47)$
IGCI $(0.06\pm 0.06)$ IGCI $(0.06\pm 0.06)$ IGCI $(0.06\pm 0.06)$ IGCI $(0.06\pm 0.06)$
PWGC $(0.05\pm 0.02)$ PWGC $(0.05\pm 0.02)$ PWGC $(0.05\pm 0.02)$ PWGC $(0.05\pm 0.02)$
RECI $(0.04\pm 0.01)$ RECI $(0.04\pm 0.01)$ RECI $(0.04\pm 0.01)$ RECI $(0.04\pm 0.01)$
Instantaneous NeuralGC $(88.38\pm 35.99)$ NeuralGC $(88.10\pm 30.65)$ NeuralGC $(86.15\pm 33.63)$ NeuralGC $(88.51\pm 31.86)$
TCDF $(4.97\pm 0.15)$ TCDF $(4.92\pm 0.12)$ TCDF $(4.93\pm 0.11)$ TCDF $(4.96\pm 0.11)$
tsFCI $(0.66\pm 0.12)$ tsFCI $(0.67\pm 0.14)$ tsFCI $(0.66\pm 0.15)$ tsFCI $(0.66\pm 0.15)$
oCSE $(0.51\pm 0.29)$ oCSE $(0.49\pm 0.29)$ oCSE $(0.43\pm 0.31)$ oCSE $(0.41\pm 0.28)$
DYNOTEARS $(0.32\pm 0.14)$ DYNOTEARS $(0.28\pm 0.10)$ DYNOTEARS $(0.28\pm 0.15)$ DYNOTEARS $(0.27\pm 0.18)$
PCMCI $(0.10\pm 0.04)$ PCMCI $(0.10\pm 0.04)$ PCMCI $(0.09\pm 0.03)$ PCMCI $(0.09\pm 0.03)$
VARLiNGAM $(0.09\pm 0.10)$ VARLiNGAM $(0.09\pm 0.10)$ VARLiNGAM $(0.09\pm 0.01)$ VARLiNGAM $(0.09\pm 0.10)$
Time-delay NeuralGC $(112.34\pm 55.35)$ NeuralGC $(113.11\pm 50.86)$ NeuralGC $(105.62\pm 48.50)$ NeuralGC $(101.19\pm 42.87)$
TCDF $(4.97\pm 0.15)$ TCDF $(4.90\pm 0.15)$ TCDF $(4.85\pm 0.14)$ TCDF $(4.94\pm 0.14)$
MVGC $(1.22\pm 1.39)$ MVGC $(1.20\pm 1.37)$ MVGC $(1.24\pm 1.43)$ MVGC $(1.21\pm 1.37)$
tsFCI $(1.05\pm 0.71)$ tsFCI $(0.86\pm 0.29)$ tsFCI $(1.04\pm 0.30)$ tsFCI $(0.99\pm 0.47)$
oCSE $(0.60\pm 0.42)$ oCSE $(0.59\pm 0.38)$ oCSE $(0.58\pm 0.35)$ oCSE $(0.56\pm 0.39)$
PCMCI $(0.52\pm 0.37)$ PCMCI $(0.46\pm 0.21)$ PCMCI $(0.49\pm 0.24)$ PCMCI $(0.53\pm 0.36)$
VARLiNGAM $(0.13\pm 0.14)$ VARLiNGAM $(0.13\pm 0.14)$ DYNOTEARS $(0.16\pm 0.16)$ DYNOTEARS $(0.18\pm 0.16)$
DYNOTEARS $(0.08\pm 0.08)$ DYNOTEARS $(0.06\pm 0.05)$ VARLiNGAM $(0.13\pm 0.14)$ VARLiNGAM $(0.13\pm 0.14)$
i.i.d GraNDAG $(211.26\pm 5.44)$ GraNDAG $(213.19\pm 19.10)$ CORL $(248.73\pm 31.10)$ CORL $(237.54\pm 1.79)$
CORL $(194.95\pm 4.56)$ CORL $(186.32\pm 1.91)$ GraNDAG $(224.22\pm 28.42)$ GraNDAG $(214.67\pm 2.71)$
DAG-GNN $(92.73\pm 102.88)$ RCD $(115.50\pm 320.54)$ DAG-GNN $(95.26\pm 98.78)$ DAG-GNN $(95.89\pm 106.63)$
GOLEM $(28.24\pm 2.61)$ DAG-GNN $(97.33\pm 112.08)$ GOLEM $(27.69\pm 3.23)$ GOLEM $(26.79\pm 1.62)$
NOTEARS $(3.20\pm 1.62)$ GOLEM $(26.57\pm 1.69)$ RCD $(4.24\pm 10.83)$ RCD $(11.68\pm 24.02)$
ES $(0.61\pm 0.27)$ NOTEARS $(4.34\pm 4.49)$ NOTEARS $(0.66\pm 0.28)$ NOTEARS $(0.70\pm 0.47)$
GES $(0.54\pm 0.16)$ ES $(0.70\pm 0.33)$ GES $(0.39\pm 0.17)$ GES $(0.33\pm 0.17)$
ICALiNGAM $(0.30\pm 0.16)$ GES $(0.49\pm 0.16)$ ES $(0.35\pm 0.21)$ ES $(0.29\pm 0.20)$
DirectLiNGAM $(0.10\pm 0.05)$ DirectLiNGAM $(0.10\pm 0.05)$ ICALiNGAM $(0.22\pm 0.09)$ ICALiNGAM $(0.28\pm 0.13)$
GRaSP $(0.09\pm 0.04)$ GRaSP $(0.10\pm 0.05)$ DirectLiNGAM $(0.11\pm 0.05)$ DirectLiNGAM $(0.10\pm 0.05)$
PC $(0.09\pm 0.05)$ PC $(0.09\pm 0.07)$ PC $(0.05\pm 0.04)$ PC $(0.06\pm 0.08)$
RCD $(0.08\pm 0.07)$ FCI $(0.08\pm 0.02)$ FCI $(0.05\pm 0.03)$ FCI $(0.06\pm 0.08)$
FCI $(0.07\pm 0.02)$ ICALiNGAM $(0.04\pm 0.03)$ GRaSP $(0.04\pm 0.03)$ GRaSP $(0.04\pm 0.04)$
Considering time cost of pairwise algorithms, RECI, PWGC, and IGCI stand out as the exemplars of efficiency across all datasets, showcasing a runtime nearly one order of magnitude lower than that of other algorithms.
For instantaneous and time-delay causality, DYNOTEARS, PCMCI, VARLiNGAM algorithm are preferred. In contrast, the NeuralGC algorithm exhibits the most prolonged computational execution time, exceeding that of the best algorithm by nearly three orders of magnitude.
When considering i.i.d. data, FCI, ICALiNGAM, and GRaSP are the preferred choices. In contrast, the least efficient algorithms, GraNDAG and CORL, have runtime that are thousands of times longer than the most efficient ones.
6.1.6 Recommendation Algorithms
It is important to note that the ranking table for each data category is obtained by calculating the average value of the results run on 15 datasets. To answer RQ 1, by sorting out the experimental findings of these causal discovery algorithms, we select the top three with the best average value under each evaluation metric as our recommended algorithms, as shown in Table 10.
Table 10: Recommendation algorithms table
F1 score AUROC FPR SHD Run Time (s)
Pairwise linear CCM CCM ANM CCM RECI
PAI PAI (L) CCM PAI PWGC
PWGC CDS CDS (S) PWGC IGCI
nonlinear PAI PAI PAI PAI RECI
CDS CDS CDS CDS PWGC
CCM CCM ANM CCM IGCI
Instantaneous linear + Gaussian oCSE oCSE oCSE oCSE VARLiNGAM
PCMCI PCMCI PCMCI PCMCI PCMCI
VARLiNGAM NeuralGC tsFCI (S) VARLiNGAM (S) DYNOTEARS
linear + non-Gaussian VARLiNGAM VARLiNGAM VARLiNGAM VARLiNGAM VARLiNGAM
oCSE oCSE oCSE oCSE PCMCI
PCMCI PCMCI PCMCI PCMCI DYNOTEARS
nonlinear + Gaussian NeuralGC NeuralGC oCSE PCMCI VARLiNGAM
tsFCI (S) oCSE PCMCI oCSE PCMCI
PCMCI (L) PCMCI tsFCI VARLiNGAM (S) DYNOTEARS
nonlinear + non-Gaussian NeuralGC NeuralGC oCSE oCSE VARLiNGAM
VARLiNGAM (S) tsFCI (L) PCMCI VARLiNGAM PCMCI
tsFCI oCSE tsFCI PCMCI DYNOTEARS
Time-delay linear + Gaussian VARLiNGAM (L) MVGC MVGC VARLiNGAM (L) DYNOTEARS
MVGC (S) VARLiNGAM VARLiNGAM MVGC VARLiNGAM
DYNOTEARS DYNOTEARS TCDF (L) PCMCI (S) PCMCI
oCSE (S)
linear + non-Gaussian PCMCI MVGC MVGC VARLiNGAM DYNOTEARS
VARLiNGAM VARLiNGAM VARLiNGAM PCMCI VARLiNGAM
DYNOTEARS PCMCI TCDF (L) MVGC (S) PCMCI
oCSE (S)
nonlinear + Gaussian PCMCI PCMCI MVGC PCMCI VARLiNGAM
DYNOTEARS MVGC VARLiNGAM (L) DYNOTEARS DYNOTEARS
VARLiNGAM DYNOTEARS oCSE (S) MVGC (S) PCMCI
nonlinear + non-Gaussian PCMCI PCMCI MVGC PCMCI VARLiNGAM
NeuralGC MVGC oCSE (S) MVGC DYNOTEARS
DYNOTEARS (S) NeuralGC VARLiNGAM (L) oCSE PCMCI
NeuralGC (L)
i.i.d. linear + Gaussian GOLEM GOLEM GOLEM GOLEM FCI
CORL CORL CORL CORL RCD
DAG-GNN DAG-GNN NOTEARS DAG-GNN PC
DAG-GNN (L) DAG-GNN (L)
linear + non-Gaussian DirectLiNGAM DirectLiNGAM DirectLiNGAM DirectLiNGAM ICALiNGAM
CORL CORL CORL CORL FCI
GOLEM GOLEM ICALiNGAM ICALiNGAM PC
nonlinear + Gaussian GraNDAG (L) CORL NOTEARS CORL GRaSP
CORL NOTEARS DAG-GNN ICALiNGAM FCI
GES GraNDAG (L) ICALiNGAM NOTEARS PC
GRaNDAG (L) GRaNDAG (L)
nonlinear + non-Gaussian CORL CORL NOTEARS CORL GRaSP
GOLEM GOLEM DAG-GNN GOLEM FCI
GraNDAG DAG-GNN ICALiNGAM ICALiNGAM PC
Note that if the optimal algorithm for a specific data size is not among the top three in average ranking, we will supplement it with gray font to ensure that Table 10 covers all possible scenarios as much as possible. “S” indicates that the algorithm is more suitable on small sample length ($L<1000$), while “L” indicates that the algorithm is more suitable on large sample length ($L>1000$).
6.2 Answer to RQ2: Real-World Applicability
We first tested the real-world dataset, Tuebingen, which comprises 100 pairs of causal relationships within a stationary time series framework characterized by nonlinearity and non-Gaussian attributes. The time length spans from 94 to 16,382 time points. Leveraging the insights posited in Section 6.1, it is deduced that the PAI, CDS, CCM algorithms are the preeminent choices under metrics such as F1 score, AUROC, and SHD; PAI, CDS, ANM are recommended for FPR metric. Additionally, RECI, PWGC, IGCI algorithm are identified as the most efficient. The dataset consists of one hundred instances of causal pairs, which we divided based on their temporal extent: those exceeding 1000 time points were classified as “large datasets” and those below 1000 time points were categorized as “small datasets”. Subsequently, an exhaustive execution of all algorithms on this real dataset was conducted, resulting in Figure 9.
Figure 9: Evaluation of algorithms (ANM, IGCI, CCM, PAI, CDS, RECI, PWGC) on real datasets: Tuebingen
Observation of the graph reveals a clear pattern: using the PAI algorithm as the benchmark, except for CDS, the average F1 and AUROC metrics of the other algorithms consistently reside beneath the horizontal baseline, while the metrics of FPR and SHD exhibit values surpassing those of PAI. This collective trend signifies that PAI demonstrates superiority as the optimal algorithm across these four evaluative metrics on “small datasets”, while CDS performs better on “large datasets”.
Moreover, when temporal considerations are factored in, the violin plot corresponding to PECI, IGCI, PWGC, CCM algorithms are prominently positioned beneath the horizontal reference line. This distinctive placement underscores that RECI holds the lowest time complexity.
These findings align with the algorithmic recommendations derived from experiments on the authentic dataset and corroborate the deductions drawn based on the outcomes expounded in Section 6.3. This congruence augments our confidence in extending the theoretical framework to real-world datasets, thereby validating our theoretical assertions and demonstrating their practical applicability.
The second real dataset pertains to functional Magnetic Resonance Imaging (fMRI), comprising 28 sets of multivariate time series. A subset of data that failed to meet criteria associated with causal sufficiency was omitted, resulting in the examination of 27 datasets. This dataset is characterized by nonlinearity and Gaussian attributes, emblematic of time-delay causal causality with $lag=1$. Among the dataset constituents, 21 sets comprise fewer than 1000 data points, while the remaining six sets exceed this threshold. Each dataset includes 5, 10, or 15 time series variables. Guided by these salient attributes, we predict that one of PCMCI, DYNOTEARS, VARLiNGAM algorithms will exhibit optimal performance under F1 score, whereas PCMCI, MVGC, DYNOTEARS algorithms will attain primacy in terms of AUROC. For FPR, MVGC is the most recommended algorithm since it performs well on datasets of all sizes, while VARLiNGAM is only recommended on large sample sizes and oCSE is only recommended on small sample sizes. Considering SHD, PCMCI, DYNOTEARS, and MVGC are recommended algorithms, with MVGC being more suitable for small sample sizes. When taking into account the time cost, we recommend PCMCI, VARLiNGAM, and DYNOTEARS algorithms.
Figure 10: Evaluation of algorithms (PCMCI, TCDF, VARLiNGAM, oCSE, MVGC, NeuralGC, DYNOTEARS, tsFCI) on real datasets: fMRI
To ensure a comprehensive assessment of algorithmic efficacy, we evaluated both instantaneous causal discovery algorithms and time-delay algorithms on the fMRI dataset. A total of nine distinct algorithms were compared. Analysis of Figure 10 reveals that the reference algorithm, PCMCI, achieves the highest values under the F1 metric and the lowest values for Run time. Regarding AUROC, FPR and SHD, it can be clearly seen from the graph that MVGC is the best performing one, which is included in our recommendations. This means that MVGC has good stability and low error rate. This alignment with our earlier assessments further bolsters the robustness of our conclusions.
To answer RQ 2, the optimal algorithm identified through experimental testing on the two datasets aligns with the one determined according to the insights in Section 6.1. This consistency underscores the reliability of the theory derived from both synthetic and real-world datasets.
6.3 Answer to RQ3: Generalization to Unknown Datasets
In Section 6.3.1, a metadata detection program was designed to extract data features. Subsequently, in Section 6.3.2, our recommendation program was tested on various datasets to verify its consistency with the algorithm test results.
6.3.1 Answer to RQ 3.1: Metadata Recognition for Algorithm Selection
Given that our prior analyses focused on causality types, linearity among series, and noise distribution, capturing these pivotal attributes within an unknown dataset is crucial for the project’s universality and practical applicability.
The first task of metadata detection is identifying temporal lags within variables. We employ the Time Lag Cross Correlation (TLCC) technique to accomplish this. TLCC is measured by gradually shifting a time series vector and repeatedly calculating the correlation between two signals. Identifying correlation maxima facilitates the ascertainment of inter-variable temporal lag. Specifically, a zero lag denotes an instantaneous causal association, whereas a non-zero lag signifies a time-delay causality. If no lag is detected, the dataset is classified as i.i.d. data.
Subsequently, identifying the noise distribution is requisite. We employ concurrent evaluative methodologies encompassing the Shapiro-Wilk, Kolmogorov-Smirnov, and Anderson-Darling tests. These tests collectively serve to discern the presence of Gaussian noise in the data. The following criteria serve as benchmarks:
1. The Shapiro-Wilk test’s computed p-value surpasses the significance threshold of 0.05.
2. The p-value resulting from the Kolmogorov-Smirnov test exceeds 0.05.
3. The p-value derived from the Anderson-Darling test remains below its critical threshold.
Fulfillment of these conditions collectively allows for the inference of Gaussian noise as the prevailing noise type characterizing the dataset.
Lastly, a crucial inquiry involves ascertaining potential linear interdependence among variables. To address this, a linear regression framework is applied to every pair of variables. Subsequently, the coefficient of determination (R-squared) is derived to gauge the efficacy of the model fit. A predetermined threshold of 0.5 is set for assessment. If the computed R-squared value surpasses this threshold, it signifies the presence of a discernible linear relationship between the variables. Conversely, an R-squared value below the threshold implies suboptimal alignment with the linear regression framework, indicating the absence of a linear relationship between the implicated variables.
To comprehensively appraise the previously delineated feature extraction procedures, we conducted metadata detection experiments on 100 datasets, with causality types, linear relations, noise distributions, and time lengths randomly generated. This evaluation was accomplished by computing the accuracy of judgments for the three metadata categories. We conducted ten trials, yielding comprehensive average and standard deviation metrics, as detailed in Table 11.
Table 11: Accuracy of metadata extraction
Causality Type Linear Relation Gaussian Noise
Accuracy $0.86\pm 0.01$ $0.81\pm 0.07$ $0.82\pm 0.09$
To answer RQ 3.1, Table 11 shows that these three metadata can be extracted precisely, with accuracies over $75\%$ and standard deviations of no more than 0.1, indicating that the judgment program is also relatively stable.
6.3.2 Answer to RQ 3.2: Practical Recommendations for Users
LUCAS and Sachs was selected as the test dataset. The extracted metadata, based on the program described in the previous section, along with the corresponding recommendation algorithms provided by Table 10, are listed in Table 12.
Table 12: Extracted metadata of LUCAS and Sachs
LUCAS Sachs
Metadata Data size (500, 11) (7466, 11)
Dependency funcs nonlinear linear
Noises distributions Gaussian non-Gaussian
Recommendation algorithms F1 GraNDAG; CORL; GES DirectLiNGAM; CORL; GOLEM
AUROC CORL; NOTEARS; GraNDAG DirectLiNGAM; CORL; GOLEM
FPR NOTEARS; DAG-GNN; ICALiNGAM DirectLiNGAM; CORL; ICALiNGAM
SHD CORL; ICALiNGAM; NOTEARS DirectLiNGAM; CORL; ICALiNGAM
Runtime GRaSP; FCI; PC ICALiNGAM; FCI; PC
13 algorithms for i.i.d. data were tested on the LUCAS dataset and 12 algorithms were tested on Sachs. The RCD algorithm is not considered for Sachs due to its long running time (run time $>15minutes$). The results, presented in Table 13 illustrates that the optimal algorithm for each metric is within previous recommendation range, which verifies the effectiveness of our recommendation program. To illustrate clearly, the algorithm we recommend is underlined, and the optimal algorithm obtained from horizontal testing is highlighted in bold.
Table 13: Test result of LUCAS and Sachs
[…]
LUCAS GraNDAG (0.70) NOTEARS (0.75) RCD (0.47) RCD (30.0) CORL (257.80)
NOTEARS (0.70) GraNDAG (0.75) DirectLiNGAM (0.33) DirectLiNGAM (16.0) GraNDAG (233.52)
GES (0.69) GES (0.72) ES (0.26) CORL (13.0) DAG-GNN (44.29)
DAG-GNN (0.69) DAG-GNN (0.72) GOLEM (0.23) ES (13.0) GOLEM (32.61)
FCI (0.67) FCI (0.68) CORL (0.23) GOLEM (11.0) RCD (12.19)
GRaSP (0.67) GRaSP (0.68) PC (0.21) PC (11.0) NOTEARS (0.83)
GOLEM (0.63) GOLEM (0.67) GES (0.16) ICALiNGAM (10.0) ES (0.71)
ICALiNGAM (0.59) ICALiNGAM (0.62) ICALiNGAM (0.16) GES (8.0) GES (0.42)
ES (0.56) ES (0.61) DAG-GNN (0.16) GRaSP (8.0) DirectLiNGAM (0.10)
CORL (0.53) CORL (0.57) GRaSP (0.09) DAG-GNN (8.0) FCI (0.05)
PC (0.48) RCD (0.52) FCI (0.09) FCI (7.0) ICALiNGAM (0.05)
RCD (0.31) PC (0.52) GraNDAG (0.02) GraNDAG (6.0) GRaSP (0.05)
DirectLiNGAM (0.21) DirectLiNGAM (0.25) NOTEARS (0.02) NOTEARS (5.0) PC (0.05)
Sachs CORL (0.30) CORL (0.35) ES (0.84) ES (37.0) DAG-GNN (814.25)
NOTEARS (0.26) DirectLiNGAM (0.32) DAG-GNN (0.70) DAG-GNN (32.0) GraNDAG (236.29)
GRaSP (0.25) GRaSP (0.32) GRaSP (0.62) FCI (31.0) CORL (95.95)
DAG-GNN (0.24) NOTEARS (0.32) PC (0.57) PC (31.0) GOLEM (68.19)
GOLEM (0.24) DAG-GNN (0.31) FCI (0.49) GRaSP (30.0) NOTEARS (25.67)
ICALiNGAM (0.23) GOLEM (0.29) CORL (0.43) CORL (25.0) ES (4.76)
PC (0.22) ICALiNGAM (0.29) NOTEARS (0.41) GES (25.0) GES (1.14)
FCI (0.19) PC (0.29) ICALiNGAM (0.35) NOTEARS (25.0) PC (1.10)
ES (0.19) ES (0.26) GOLEM (0.32) GOLEM (24.0) FCI (0.68)
DirectLiNGAM (0.17) FCI (0.26) GES (0.19) ICALiNGAM (24.0) DirectLiNGAM (0.38)
GES (0.16) GES(0.22) GraNDAG (0.14) GraNDAG (20.0) GRaSP (0.35)
GraNDAG (0.08) GraNDAG (0.18) DirectLiNGAM (0.08) DirectLiNGAM (17.0) ICALiNGAM (0.09)
To answer RQ 3.2, the case study of these two real datasets indicates that the optimal algorithms obtained from the experiment are included in our recommended algorithms. This means that users can quickly find the most suitable algorithm based on extracted metadatas and save computing power.
7 Threats to Validity
7.0.1 Threats to internal validity
(1)Correctness of The Codes. A potential threat to the study revolves around the accuracy of the employed codebase. Despite rigorous testing and validation procedures, the complexity of algorithmic implementations and potential oversights during the coding process may give rise to errors. Variations in the code may inadvertently influence the outcomes, thereby introducing a threat to the study’s internal validity.
To address this, the study highlights the ongoing commitment to code review and validation, emphasizing the significance of code quality in ensuring the reliability of the study’s outcomes. Specifically, we conducted validation tests on each algorithm to ensure its correctness.
(2) Implementation of Control Variables. In light of the comparative analytical methodology employed within the experimental framework of this project, the precision in governing variables assumes paramount importance due to its direct influence on the ensuing output. If multiple testing algorithms are administered across disparate datasets, the potential for consequential impact stemming from inherent dataset dissimilarities becomes salient. Such variations could undermine the cogency of the ultimate conclusions drawn from the study.
To mitigate this, distinct algorithms were executed on a singular dataset, thereby facilitating establishing a controlled environment wherein the dataset variable maintains consistency. Notably, our program was executed holistically on a singular server infrastructure, ensuring uniformity in the controlling variable of the computing capacity and thereby rendering temporal comparisons feasible. In the interest of upholding the validation of the comparative analysis, synthesized data types of specific time lengths were all derived from one DAG structure. This strategic alignment eradicated any potential ambiguities stemming from disparate ground truth, consequently enhancing the reliability and interpretability of the results.
7.0.2 Threats to external validity
(1) Metadata selection bias. The external validity of this study could be threatened by metadata selection bias. In light of our methodology, which exclusively constructs testing data predicated upon established causal linkages, noise distribution characteristics, and dataset dimensions, it is prudent to acknowledge the potential bias that crucial time series’ metadata might not be exhaustively contained. Consequently, there remains a prospect for more influential data attributes that could threaten the comprehensiveness and external validity of the experimental outcomes.
In order to mitigate the potential influence of this factor on the experimental validity, two real-world datasets were examined, and it was ascertained that the empirical findings obtained from these real-world datasets exhibit congruence with those derived from the synthetic datasets. This correspondence serves as an indirect validation of the judicious selection of metadata, affirming its representativeness.
(2) Temporal Validity. Considering the incessant evolution characterizing algorithms, it is plausible that our research might not encompass forthcoming optimizations, thereby engendering temporal constraints on the experiment’s conclusions. This scenario presents an external threat to the generalizability of research endeavours.
To attenuate this influence to the greatest extent possible, a mitigation strategy has been adopted wherein algorithms are encoded following a standardized input-output paradigm, thereby creating an extensible code library. This design contributes to the seamless integration of forthcoming algorithmic enhancements into the repository, facilitating their inclusion within the testing framework. This holistic program design underscores commendable operational feasibility and assimilation capacity.
8 Conclusions
This section will summarize the entire project and analyze potential future work from three aspects.
8.1 Conclusions and Summary
This work conducted comprehensive research on causal discovery in time series, introducing the research topic’s relevant background, importance, and literature system. We introduced six types of methods and sorted out over 20 algorithms.
In terms of experiments, our task is to explore the optimal algorithms in different application scenarios. We successfully summarised recommendation algorithms corresponding to 16 data types through comparative experiments and verified the applicability of these insights through testing on two real-world datasets. Furthermore, we extended the practicality of this discovery to the unknown dataset through metadata extraction technology, with the testing accuracy up to $80\%$.
Finally, we discussed the experimental results and related them to existing literature, and listed the threats that could affect the validity of the results. By discussing these two aspects, the findings of this project are made more reliable and vivid.
Given the above content, this project has systematically completed all research objectives, thoroughly analyzing causal discovery for time series from theoretical and practical aspects. We provided practical and effective algorithm recommendations from users’ perspectives, filling the research gaps in previous articles and providing specific reference values for future algorithm research.
8.2 Recommendations for Further Work
Opportunities for enhancement and progress persist in this study. Based on the discussion section, a trajectory for future investigations can be charted across three principal dimensions. The initial facet involves considering a broader range of data types, such as high-dimensional datasets or time series incorporating latent variables.
The second avenue of progression entails the extension of algorithmic testing. Due to temporal constraints in this project, certain time-intensive algorithms were omitted from the assessment. It is advisable to incorporate these additional algorithms in future endeavors to achieve a more exhaustive and comprehensive comparative analysis of algorithmic efficacy.
The final and pivotal facet pertains to enhancing the precision of metadata identification. To achieve this objective, the adoption of more sophisticated data processing techniques is warranted. Notably, the integration of machine learning methodologies should be considered. This approach would yield more precise determinations of unknown data characteristics, particularly in distinguishing between time-delay and instantaneous causality.
Upon the successful execution of the proposed endeavors, users will be furnished with a meticulous causal discovery algorithm recommendation service. This service would facilitate the expedient identification of optimal algorithms for arbitrary datasets, thereby markedly curtailing the duration of trial-and-error procedures and minimizing computational resource consumption. This work effectively addresses the intricate conundrum associated with algorithm selection for causal discovery, which holds significant research and practical value.
Additionally, future work could explore the application of causal discovery methods to verify the results of data augmentation (Gao and Li, 2024). For example, these methods could be used to ascertain whether newly generated data retains the causality of the original data. This verification process would enhance the reliability and validity of augmented datasets, ensuring that the fundamental causal relationships are preserved.
Appendix A Residual Plot
(a) Linear and Gaussian Datasets (pairwise)
(b) Linear and Non-Gaussian Datasets (pairwise)
(c) Nonlinear and Gaussian Datasets (pairwise)
(d) Nonlinear and Non-Gaussian Datasets (pairwise)
Figure A1: Evaluation of pairwise algorithms (ANM, IGCI, CCM, PAI, CDS, PECI, PWGC) on synthetic datasets. The original violin plot of the reference algorithm is presented in the left column while the corresponding residual plot of alternative algorithms is presented in the right column. The presence of an asterisk denotes statistical significance as determined by the Mann-Whitney U test with a significance threshold of $p<0.05$.
(a) Linear and Gaussian Datasets (instantaneous)
(b) Linear and Non-Gaussian Datasets (instantaneous)
(c) Nonlinear and Gaussian Datasets (instantaneous)
(d) Nonlinear and Non-Gaussian Datasets (instantaneous)
Figure A2: Evaluation of instantaneous causal algorithms (oCSE, TCDF, VARLiNGAM, NeuralGC, DYNOTEARS, PCMCI, tsFCI) on synthetic datasets. The original violin plot of the reference algorithm is presented in the left column while the corresponding residual plot of alternative algorithms is presented in the right column. The presence of an asterisk denotes statistical significance as determined by the Mann-Whitney U test with a significance threshold of $p<0.05$.
(a) Linear and Gaussian Datasets (time-delay)
(b) Linear and Non-Gaussian Datasets (time-delay)
(c) Nonlinear and Gaussian Datasets (time-delay)
(d) Nonlinear and Non-Gaussian Datasets (time-delay)
Figure A3: Evaluation of time-delay causal algorithms (MVGC, PCMCI, TCDF, VARLiNGAM, NeuralGC, oCSE, DYNOTEARS, tsFCI) on synthetic datasets. The original violin plot of the reference algorithm is presented in the left column while the corresponding residual plot of alternative algorithms is in the right column. An asterisk denotes statistical significance determined by the Mann-Whitney U test with $p<0.05$.
(a) Linear and Gaussian Datasets (i.i.d.)
(b) Linear and Non-Gaussian Datasets (i.i.d.)
(c) Nonlinear and Gaussian Datasets (i.i.d.)
(d) Nonlinear and Non-Gaussian Datasets (i.i.d.)
Figure A4: Evaluation of i.i.d. causal algorithms (RCD, GraNDAG, GES, NOTEARS, PC, GOLEM, DAG-GNN, CORL, ICALiNGAM, DirectLiNGAM, FCI, ES, GRaSP) on synthetic datasets. The original violin plot of the reference algorithm is presented in the left column while the corresponding residual plot of alternative algorithms is in the right column. An asterisk denotes statistical significance determined by the Mann-Whitney U test with $p<0.05$.